### v1

# Non-abelian Hodge correspondence over singular Kähler spaces

Abstract.
In this paper, we establish the non-abelian Hodge correspondence over compact Kähler spaces with Kawamata log terminal (klt) singularities as well as over their regular loci, thereby extending the result of Greb-Kebekus-Peternell-Taji for projective klt varieties [38, 39] to the context of compact Kähler klt spaces.
The proof relies on two key ingredients: first, we establish an equivalence over the regular loci—via harmonic bundles—between polystable Higgs bundles with vanishing orbifold Chern numbers and semi-simple flat bundles; second, we prove a descent theorem for semistable Higgs bundles with vanishing Chern classes along resolutions of singularities.
As an application of our framework, we obtain a quasi-uniformization theorem for projective klt varieties with big canonical divisor that satisfy the orbifold Miyaoka-Yau equality in the sense of [53].
Key words and phrases:
Higgs sheaves, orbifold Chern classes, non-abelian Hodge theory, Kähler spaces, Kawamata log terminal singularities, Miyaoka-Yau inequality
2020 Mathematics Subject Classification:
Primary 32J25, Secondary 32Q15, 53C07
1. Introduction
The non-abelian Hodge correspondence stems from the work of Corlette, Donaldson, Hitchin and Simpson [17, 25, 47, 97] and can be summarized by the following equivalences:
$\left\{\begin{subarray}{c}\text{polystable Higgs bundles}\\ \text{with vanishing Chern numbers}\end{subarray}\right\}$$\left\{\text{harmonic bundles}\right\}$$\left\{\begin{subarray}{c}\text{semi-simple linear representations}\\ \text{from }\pi_{1}(X)\text{ to }\mathrm{GL}(r,\mathbb{C})\end{subarray}\right\}$$\left\{\text{semi-simple flat bundles}\right\}$}Hitchin-SimpsonNon-abelian Hodge correspondenceCorlette-DonaldsonRiemann-Hilbert
In general, Simpson [99] established an equivalence between the category of semistable Higgs sheaves with vanishing Chern numbers and the category of linear representations of the fundamental group. This equivalence also holds for compact Kähler manifolds [86, 23] and has been significantly generalised to other settings, such as quasi-Kähler manifolds [98, 61, 83, 84, 85] and varieties over fields of positive characteristic [26, 87, 1, 69, 68, 96], we refer to [74, 52] for an overview of its development.
The minimal model program (MMP) for Kähler spaces has advanced substantially in recent years (see e.g. [50, 5, 6, 22, 20, 29]). A key feature is that running the MMP typically introduces singularities, even when starting from a smooth Kähler space. Consequently, establishing a singular version of the non-abelian Hodge correspondence becomes a fundamental problem.
The main results of this paper establish the non-abelian Hodge correspondence for compact Kähler spaces with Kawamata log terminal (klt) singularities, as well as over their regular loci. This work is motivated by two recent lines of progress: first, the successful construction of the correspondence for projective klt varieties [38, 39]; and second, the recent characterization of stable reflexive sheaves satisfying the orbifold Bogomolov–Gieseker inequality on compact Kähler klt spaces [90, 28] (see also [41, 42]). A further motivation comes from applications to the uniformization of Kähler klt spaces, extending analogous developments in the projective setting (cf. [39, 14, 92, 55] and references therein).
1.1. Main results
General definitions and conventions follow [110].
Notation 1.1.
Let $\mathcal{E}_{X_{\text{reg}}}$ be a coherent sheaf defined only on the regular locus of $X$. We denote by $\mathcal{E}_{X}$ its trivial extension to the whole space. Conversely, if $\mathcal{E}_{X}$ is a sheaf on $X$, we write $\mathcal{E}_{X_{\text{reg}}}$ for its restriction to the regular locus. For a morphism $f:Y\to X$ and a coherent sheaf $\mathcal{E}_{X}$ on $X$, we write $f^{[*]}\mathcal{E}_{X}:=(f^{*}\mathcal{E}_{X})^{\vee\vee}$ for the reflexive pull-back (i.e., the double dual of the pull-back).
Definition 1.2.
Let $X$ be a compact Kähler klt space of dimension $n$, we define
  • $\mathrm{Higgs}_{X}$: locally free Higgs sheaves $(\mathcal{E}_{X},\theta_{X})$ of rank $r$ on $X$ such that $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with vanishing Chern classes
    (1.1) $\mathrm{ch}_{1}(\mathcal{E}_{X})\cdot[\omega_{0}]\cdots[\omega_{n-2}]=\mathrm{ch}_{2}(\mathcal{E}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]=0$
    for some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$.
  • $\mathrm{LSys}_{X}$: local systems of rank $r$ on $X$.
Analogously, we can define the categories $\mathrm{pHiggs}_{X}$ and $\mathrm{sLSys}_{X}$ for polystable locally free Higgs sheaves satisfying (1.1) and semi-simple local systems.
The first main result of this paper is establishing the non-abelian Hodge correspondence over compact Kähler klt spaces.
Theorem 1.3.
Let $X$ be a compact Kähler klt space. There exists a natural one-to-one correspondence
$\mu_{X}:\mathrm{Higgs}_{X}\longrightarrow\mathrm{LSys}_{X}$
which is compatible with resolutions of singularities in the following sense. For any resolution $\pi:\widehat{X}\to X$ and any $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$, the pullback $\pi^{*}(\mathcal{E}_{X},\theta_{X})$ belongs to $\mathrm{Higgs}_{\widehat{X}}$ and the diagram
(1.2) $\mathrm{Higgs}_{X}$$\mathrm{LSys}_{X}$$^{\mathrm{Higgs}_{\widehat{X}}}$$^{\mathrm{LSys}_{\widehat{X}}}$$\scriptstyle{\mu_{X}}$$\scriptstyle{\pi^{*}}$$\scriptstyle{\pi^{*}}$$\scriptstyle{\mu_{\widehat{X}}}$
commutes. Here $\mu_{\widehat{X}}$ denotes the classical non-abelian Hodge correspondence for the compact Kähler manifold $\widehat{X}$.
Remark 1.4.
We refer to Remark 2.12 for the definition of the pullback $\pi^{*}(\mathcal{E}_{X},\theta_{X})$. Moreover, the restriction of $\mu_{X}$ to $\mathrm{pHiggs}_{X}$ yields a one-to-one correspondence from $\mathrm{pHiggs}_{X}$ to $\mathrm{sLSys}_{X}$ that is likewise compatible with any resolution of singularities.
It is necessary to consider a distinct correspondence $\mu_{X_{\mathrm{reg}}}$ over the regular locus of a compact Kähler klt space $X$, because the induced homomorphism $i_{*}:\pi_{1}(X_{\mathrm{reg}})\to\pi_{1}(X)$ is surjective (cf. [65, Proposition 2.10]) but generally not injective.
Definition 1.5.
Let $X$ be a compact Kähler klt space $X$ of dimension $n$, we define
  • $\mathrm{Higgs}_{X_{\mathrm{reg}}}$: reflexive Higgs sheaves $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ of rank $r$ on $X_{\mathrm{reg}}$ such that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with vanishing orbifold Chern classes
    (1.3) $\widehat{\mathrm{ch}}_{1}(\mathcal{E}_{X})\cdot[\omega_{0}]\cdots[\omega_{n-2}]=\widehat{\mathrm{ch}}_{2}(\mathcal{E}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]=0$
    for some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$.
  • $\mathrm{LSys}_{X_{\mathrm{reg}}}$: local systems of rank $r$ on $X$.
Analogously, we can define the categories $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ and $\mathrm{sLSys}_{X_{\mathrm{reg}}}$ for polystable reflexive Higgs sheaves satisfying (1.3) and semisimple local systems.
Remark 1.6.
Our definitions employ the orbifold Chern classes, denoted $\widehat{\mathrm{ch}}_{1}$ and $\widehat{\mathrm{ch}}_{2}$, which coincide with $\mathbb{Q}$-Chern classes upon intersection with ample divisors. In particular, the definition of $\mathrm{Higgs}_{X_{\mathrm{reg}}}$ coincides with [39] when $\omega_{0}],\cdots,[\omega_{n-2}]$ are ample. We refer to [90, Section 2] for an introduction of the precise definitions of orbifold Chern classes.
The second main result of this paper is establishing the non-abelian Hodge correspondence over the regular loci of compact Kähler klt spaces.
Theorem 1.7.
Let $X$ be a compact Kähler klt space. There exists a one-to-one correspondence
$\mu_{X_{\mathrm{reg}}}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\longrightarrow\mathrm{LSys}_{X_{\mathrm{reg}}},$
such that the underlying $C^{\infty}$-bundle of $\mathcal{E}_{X_{\mathrm{reg}}}$ admits a flat connection induced by $\rho_{X_{\mathrm{reg}}}$. Moreover, this correspondence is compatible with the correspondence from Theorem 1.3 in the following sense: for any maximally quasi-étale cover $\gamma:Y\to X$, the reflexive pullback $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})$ belonging to $\mathrm{Higgs}_{Y}$, and the diagram
(1.4) $\mathrm{Higgs}_{X_{\mathrm{reg}}}$$\mathrm{LSys}_{X_{\mathrm{reg}}}$$\mathrm{Higgs}_{Y}$$\mathrm{LSys}_{Y}$$\scriptstyle{\mu_{X_{\mathrm{reg}}}}$$\scriptstyle{\gamma^{*}}$$\scriptstyle{\gamma^{*}}$$\scriptstyle{\mu_{Y}}$
commutes.
Remark 1.8.
Here, $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ denotes the reflexive pullback (cf. Lemma 2.10).
  1. (i)
    As in [38, 39], maximally quasi-étale covers take a central role in the construction of $\mu_{X}$ and $\mu_{X_{\mathrm{reg}}}$. We refer to Section 4.1 for its definitions.
  2. (ii)
    The restriction of $\mu_{X_{\mathrm{reg}}}$ to $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ gives the bijection from $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ to $\mathrm{sLSys}_{X_{\mathrm{reg}}}$ that is compatible with every maximally quasi-étale cover.
Together, Theorems 1.3 and 1.7 provide the analytic counterpart of the two correspondences established for projective klt varieties in [38] and [39], respectively.
1.2. Applications to quasi-uniformization problems
We now present several applications of our main results. First, we give a numerical characterization of singular quotients of complex tori, established via different methods in [14, 42] for a single Kähler form (cf. [75, 35, 13] for previous advances). The smooth version of this result originates in Yau’s solution of the Calabi conjecture [109].
Corollary 1.9.
Let $X$ be a compact Kähler klt space of dimension $n$ with $K_{X}$ numerically trivial. Then
(1.5) $\widehat{c}_{2}(X)\cdot[\omega_{1}]\cdots[\omega_{n-2}]\geq 0,$
with equality if and only if $X$ is a singular quotient of a torus $\mathbb{T}^{n}$; i.e., there exists a finite group $G\subset\operatorname{Aut}(\mathbb{T}^{n})$ acting freely in codimension one such that $X\cong\mathbb{T}^{n}/G$.
When $K_{X}$ is nef, inequality (1.5)—known as Miyaoka’s inequality, originating from [82]—continues to hold in the Kähler klt setting (cf. [53, Theorem 1.5]). Its equality case has been characterized for compact Kähler manifolds [88, 54, 79] and for projective klt varieties [55]. Extending these results to compact Kähler klt spaces requires a characterization of the semistable reflexive Higgs sheaves that satisfy the orbifold Bogomolov–Gieseker equality [53, Question 7.5], which has been proved for stable reflexive Higgs sheaves in [28] (see also [41, 42]). We now provide such a characterization based on Theorems 1.3 and 1.7, combined with the preservation of semistability under tensor products proved in [110].
Corollary 1.10.
Let $X$ be a compact Kähler klt space and let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover. Suppose $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a reflexive Higgs sheaf of rank $r$ on the regular locus, semistable with respect to some Kähler forms $\omega_{0},\dots,\omega_{n-2}$. Then the following orbifold Bogomolov–Gieseker inequality holds:
(1.6) $\Bigl(2\widehat{c}_{2}(\mathcal{E}_{X})-\frac{r-1}{r}\widehat{c}_{1}^{2}(\mathcal{E}_{X})\Bigr)\cdot[\omega_{1}]\cdots[\omega_{n-2}]\geq 0.$
Equality in (1.6) holds if and only if the sheaf $\mathcal{G}_{X}:=\gamma^{[*]}\operatorname{End}(\mathcal{E}_{X})$ is locally free and the induced Higgs sheaf $(\mathcal{G}_{X},\theta_{\mathcal{G}})$ corresponds, via the correspondence $\mu_{X}$ of Theorem 1.3, to a linear representation $\rho:\pi_{1}(Y)\to\operatorname{GL}(r,\mathbb{C})$.
As another key application, we prove a quasi‑uniformization theorem for projective klt varieties with big canonical divisor. The Miyaoka–Yau inequality, originating in [81, 82], is a cornerstone in the geometry of surfaces of general type; see the introductions of [37, 43] for an overview. For a minimal projective klt variety $X$ of general type, the authors of [37] established the orbifold Miyaoka-Yau inequality
(1.7) $\bigl(2\widehat{c}_{2}(X)-\tfrac{n}{n+1}\widehat{c}_{1}^{2}(X)\bigr)\cdot K_{X}^{n-2}\geq 0.$
This has been extended to minimal projective dlt pairs [43] and, very recently, to minimal Kähler klt spaces [110] (see also [53]). Moreover, the equality case of (1.7) yields a quasi‑uniformization description [39]: the canonical model of $X$ is a singular quotient of the unit ball.
Making use of the framework developed in this paper, we now extend this picture to varieties whose canonical divisor is big, but not necessarily nef.
Theorem 1.11.
Let $X$ be a projective klt variety of dimension $n$ with $K_{X}$ big. The following orbifold Miyaoka-Yau type inequality holds:
(1.8) $\bigl(2\widehat{c}_{2}(X)-\tfrac{n}{n+1}\widehat{c}_{1}^{2}(X)\bigr)\cdot\langle K_{X}^{n-2}\rangle\geq 0.$
If equality holds, then the canonical model $X_{\mathrm{can}}$ of $X$ is a singular quotient of a projective manifold $Y$ that is covered by the unit ball $\mathbb{B}^{n}$.
The inequality (1.8) was established in the recent work [53]—a result that is new even in the smooth case. Here $\langle K_{X}^{n-2}\rangle$ denotes the non-pluripolar product; the precise definition of the intersection number appearing in (1.8) is given in [53, Section 4.1]. Together with the very recent preprint [58]—which proves that for a $K$-stable projective klt variety with big anti-canonical divisor, the Miyaoka-Yau type equality forces its canonical model to be a singular quotient of $\mathbb{P}^{n}$—these results confirm that the inequality (1.8) developed in [53] provides an effective criterion for projective klt varieties with big canonical or anti-canonical divisors.
1.3. Overview and main technical components
We now outline the proofs of Theorems 1.3 and 1.7 and present the principal differences from the projective case treated in [38, 39].
Let $X$ be a compact Kähler klt space. Following the strategy of [38], a key step in constructing $\mu_{X}$ is to show that for any $(\mathcal{E}_{X},\theta_{X})\in\operatorname{Higgs}_{X}$ and any resolution $\pi:\widehat{X}\to X$, the pullback $\pi^{*}(\mathcal{E}_{X},\theta_{X})$ lies in $\operatorname{Higgs}_{\widehat{X}}$. In Section 5.1 we prove the following.
Theorem 1.12 (cf. Theorems 5.4 and 5.5).
Let $X$ be a compact Kähler klt space and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ a reflexive Higgs sheaf on $X_{\mathrm{reg}}$.
  • (1)
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$ (resp. $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$) if and only if for every maximally quasi-étale cover $\gamma:Y\rightarrow X$, $\gamma^{[*]}\mathcal{E}_{X}$ is locally free and $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{Higgs}_{Y}$ (resp. $\mathrm{pHiggs}_{Y}$). Moreover, $(\mathcal{E}_{Y},\theta_{Y})$ is an extension of locally free Higgs sheaves belonging to $\mathrm{pHiggs}_{Y}$.
  • (2)
    If $\mathcal{E}_{X}$ is locally free, then $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ (resp. $\mathrm{pHiggs}_{X}$) if and only if $\pi^{*}(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{\widehat{X}}$ (resp. $\mathrm{pHiggs}_{\widehat{X}}$) for every resolution of singularities $\pi:\widehat{X}\rightarrow X$.
The main task is to prove that for any Jordan–Hölder filtration $\{\mathcal{E}_{i}\}_{i=0}^{l}$ of $\gamma^{[*]}(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})$, each quotient sheaf $\mathcal{E}_{i}/\mathcal{E}_{i-1}$ is locally free and has vanishing Chern classes. Following the spirits of [38, Proposition 7.10], [48, Theorem 1.2] and [55, Proposition 3.10], the problem in our context reduces to proving the following statement: if $(\mathcal{Q}_{Y_{\mathrm{reg}}},\theta_{\mathcal{Q}_{Y_{\mathrm{reg}}}})\in\operatorname{pHiggs}_{Y_{\mathrm{reg}}}$ is stable, then its reflexive extension $\mathcal{Q}_{Y}$ is locally free and all its Chern classes vanish. In the projective case [38, Proposition 7.10], this is handled by cutting down with hypersurface sections to reduce to surfaces, where reflexive sheaves are automatically locally free—a technique unavailable for general Kähler spaces. The central difficulty in the Kähler setting is therefore precisely to establish the local freeness of $\mathcal{Q}_{X}$ in higher dimensions.
In Sections 3 and 4, we concentrate on proving the following characterization of polystable reflexive Higgs sheaves satisfying (1.3). This result not only yields the desired local freeness of $\mathcal{Q}_{Y}$ but simultaneously constructs the correspondence $\mu_{X_{\mathrm{reg}}}:\mathrm{pHiggs}_{X_{\mathrm{reg}}}\to\mathrm{sRep}_{X_{\mathrm{reg}}}$.
Theorem 1.13.
Let $X$ be a compact Kähler klt space and $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover.
  • (1)
    A flat bundle $(E_{X_{\mathrm{reg}}},D_{X_{\mathrm{reg}}})$ is semi-simple if and only if it admits a pluri-harmonic metric $H_{X_{\mathrm{reg}}}$, which is unique up to a positive constant on each simple component.
  • (2)
    A reflexive Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ belongs to $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ if and only if $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ arises from a harmonic bundle $(E_{X_{\mathrm{reg}}},D_{X_{\mathrm{reg}}},H_{X_{\mathrm{reg}}})$; here $H_{X_{\mathrm{reg}}}$ is unique up to obvious ambiguity. This is also equivalent to the one of the following three statements:
    1. (i)
      The reflexive pullback $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{pHiggs}_{Y}$ with vanishing Chern classes.
    2. (ii)
      For any resolution of singularities $\pi:\widehat{Y}\rightarrow Y$, $\pi^{*}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{pHiggs}_{\widehat{Y}}$.
    3. (iii)
      $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is polystable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ and all orbifold Chern classes of $\mathcal{E}_{X}$ vanish.
The proof of Theorem 1.13 addresses this issue and proceeds in three steps.
  • Step 1.
    In Section 3, we prove the existence of a pluriharmonic metric for reflexive Higgs sheaves, which had been proceeded in [28] for the case where $\theta_{X_{\mathrm{reg}}}$ is trivial (see also [41, 42]). Our proof adapts the Higgs version [73] of Bando-Siu’s argument [3] to study the Hermitian-Yang-Mills flow $\{H_{\mathrm{orb}}(t)\}$ on a partial orbifold modification constructed in [67]. Combining this with (1.3), we conclude that the limiting Hermitian–Einstein metric is pluri-harmonic. The overall strategy are detailed in Section 3.1.
  • Step 2.
    In the first part of Section 4.2, we prove part (1) of the theorem. The existence of pluri-harmonic metrics on semi-simple flat bundles over the regular locus follows from the work of Corlette and Donaldson [17, 25] via maximally quasi‑étale covers; the remaining properties are obtained through the maximum principle, which is well‑adapted to our setting thanks to the extension theorem of Grauert and Remmert [32].
  • Step 3.
    From Step 1 and Theorem 1.13 (1), we conclude that $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ admits a pluri-harmonic metric induced by a semi-simple local system $E_{Y_{\mathrm{reg}}}$. Since $Y$ is maximally quasi-étale, $E_{Y_{\mathrm{reg}}}$ extends to a semi-simple local system $E_{Y}$. To verify that $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf on $Y$, it suffices to establish a descent result for Higgs bundles arising from harmonic bundles (Proposition 4.11), which we prove using the period map constructed in [19]. Details are given in Sections 4.2 and 4.3.
Remark 1.14.
The careful treatment of harmonic bundles on the regular locus presented above is necessary in the context of compact Kähler klt spaces because Mochizuki’s theory of tame harmonic bundles on quasi-projective manifolds [83, 84, 85]—which crucially underlies the construction of $\mu_{X_{\mathrm{reg}}}$ in the projective case [39]—has not yet been established for general quasi-Kähler manifolds.
It should be noted that the admissibility of the limiting Hermitian–Einstein metric constructed in Section 3 remains unknown (cf. Question 3.6). The key point of Steps 2 and 3 is that klt singularities are sufficiently mild to allow the characterization of polystable Higgs bundles satisfying (1.3), or of semisimple local systems via harmonic bundles on the regular locus, without requiring any extra assumptions on the pluriharmonic metrics.
Building on Theorem 3.2, we confirm that each $(\mathcal{E}_{i}/\mathcal{E}_{i-1})^{\vee\vee}$ is locally free. Then, in Section 5.1, we complete the proof of Theorem 1.12. This allows us to construct $\mu_{X}$ and $\mu_{X_{\mathrm{reg}}}$ in Section 5.3. It remains to prove that both maps are injective; this relies on proving the following descent result as in [38].
Theorem 1.15.
Let $f:Z\rightarrow X$ be a bimeromorphic holomorphic map between compact Kähler klt spaces. Then for any $(\mathcal{E}_{Z},\theta_{Z})\in\mathrm{Higgs}_{Z}$, there exists $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ such that $(\mathcal{E}_{Z},\theta_{Z})=f^{*}(\mathcal{E}_{X},\theta_{X})$.
Theorem 1.15 was proved for locally free $f$-numerically flat sheaves in the projective case [39, Theorem 1.2], building on the relative MMP for projective morphisms [4]. Its extension to the Kähler klt setting seems plausible, especially in light of recent work extending the relative MMP to analytic geometry [20, 29]. However, several technical results used in the proof of [38]—notably [49, Proposition 3.1], which appears in [38, Claim 4.5]—are, to the best of our knowledge, not directly available in the literature for the analytic setting. Moreover, in order to maintain our focus on Higgs sheaves and to keep the paper concise, we develop in Section 5.2 an alternative argument. It factors through local systems via an orbifold modification, uses the Jordan–Hölder filtration of the corresponding representation, and reconstructs the Higgs sheaf by descending the graded stable pieces based on Proposition 4.11.
Our overarching goal—proving Theorems 1.12 and 1.15—is the same as in [38, 39], the proofs we present in the Kähler setting are substantially different, for the reasons outlined above.
Acknowledgement. The authors wish to thank Xin Fu for useful discussions on uniform geometric estimates; Changpeng Pan and Di Wu for valuable conversations about harmonic bundles; and Masataka Iwai and Satoshi Jinnouchi for sharing their perspectives on quasi-uniformization problems. The second author acknowledges Tianshu Jiang for communications on related topics. The research was supported by the National Key R and D Program of China 2020YFA0713100. The authors are partially supported by NSF in China No.12141104, 12371062 and 12431004.
2. Fundamental materials
This section reviews and proves necessary statements concerning complex orbifolds (Section 2.1), Higgs sheaves (Section 2.2), orbifold Chern classes (Section 2.3) and quasi-étale morphisms (Section 2.4).
2.1. Complex orbifolds
The study on orbifolds originated from [94]. For a concise discussion to complex orbifolds, we refer to [22, Section 3] and [107].
A complex space $X$ with quotient singularities equipped with an effective orbifold structure $X_{\mathrm{orb}}:=\{V_{\alpha},G_{\alpha},\mu_{\alpha}\}$, namely, there exists an open cover $\{X_{\alpha}\}$ of $X$ such that $X_{\alpha}\cong V_{\alpha}/G_{\alpha}$ for some complex manifold $V_{\alpha}$ and $G_{\alpha}\subset\mathrm{Aut}(V_{\alpha})$ and $(V_{\alpha},G_{\alpha},\mu_{\alpha})$ are compatible along the overlaps, where $\mu_{\alpha}:V_{\alpha}\rightarrow X_{\alpha}$ is the quotient map induced by $G_{\alpha}$. Effective means that every $G_{\alpha}$-action is faithful. Indeed, $X$ always admits a standard orbifold structure, i.e. $\pi_{\alpha}$ is étale in codimension $1$. In virtue of puriry theorem of branch locus due to Grauert-Remmert [32], it can be easily seen that
Lemma 2.1 (c.f. [107, Proposition 1]).
Let $\{V_{\alpha},G_{\alpha},\pi_{\alpha}\}$ and $\{V_{\beta},G_{\beta},\pi_{\beta}\}$ be two standard orbifold structure. Then $V_{\alpha,\beta}:=\{(z_{\alpha},z_{\beta})\in V_{\alpha}\times V_{\beta}:\mu_{\alpha}(z_{\alpha})=\mu_{\beta}(z_{\beta})\}$ is smooth and the the natural projection $V_{\alpha,\beta}\rightarrow V_{\alpha}$ and $V_{\alpha,\beta}\rightarrow V_{\beta}$ are étale.
By definition, $V_{\alpha,\beta}/(G_{\alpha}\times G_{\beta})\cong X_{\alpha}\cap X_{\beta}$. Then by Zorn’s Lemma, there exists a standard orbifold structure that is maximal. When we consider a standard orbifold structure, we will always refer to the maximal one.
Lemma 2.2 (cf. [22, Lemma 3.10]).
Let $W_{orb}=(V_{i},G_{i})$ be a compact complex orbifold, and $\mathcal{E}_{orb}=\{\mathcal{E}_{i}\}$ be an torsion-free orbi-sheaf on $W_{orb}$. Then there exists an orbifold morphism $h_{orb}:Z_{orb}\rightarrow W_{orb}$ from a compact complex orbifold $Z_{orb}=\{(\widehat{V}_{i},G_{i},\widehat{\pi}_{i})\}$ so that the following holds.
  • (1)
    Set $h_{orb}:=\{h_{i}:\widehat{V}_{i}\rightarrow V_{i}\}$, then $h_{i}$ is obtained by a sequence of blowups of $G_{i}$-invariant smooth centers contained in the non-locally-free locus of $\mathcal{E}_{i}$.
  • (2)
    The coherent orbi-sheaf $E_{\mathrm{orb}}:=\{h_{i}^{*}\mathcal{E}_{i}/(\mathrm{torsion})\}$ is a vector orbi-bundle.
  • (3)
    If $W_{orb}$ is Kähler, then $Z_{orb}$ is also Kähler.
All objects (e.g. orbi sheaves, orbifold differential forms, etc.,) in the context of orbifolds are those defined on each $V_{\alpha}$ satisfying compatibility conditions. For a concise introduction to Higgs orbi-sheaves, we refer to [110, Section 2].
2.2. Higgs sheaves
We refer to [110, Section 4] for a concise discussion on Higgs sheaves over the regular locus of a compact complex normal space $X$. All discussions are still valid for locally free Higgs sheaves over $X$ by the following Remark 2.8.
2.2.1. Higgs sheaves over the whole space
We follow [37, Section 5] to introduce Higgs sheaves over normal spaces (see also [53, Section 3.2.2]).
Definition 2.3 (Higgs sheaves, cf. [37, Definition 5.1]).
Let $X$ be a normal space. A Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ on $X$ consists of a coherent sheaf and a morphism $\theta_{X}:\mathcal{E}_{X}\rightarrow\mathcal{E}_{X}\otimes\Omega_{X}^{[1]}$ such that the composed morphism
$\theta\wedge\theta:\mathcal{E}_{X}\xrightarrow{\theta}\mathcal{E}_{X}\otimes\Omega_{X}^{[1]}\xrightarrow{\theta\otimes\operatorname{Id}}\mathcal{E}_{X}\otimes\Omega_{X}^{[1]}\otimes\Omega_{X}^{[1]}\xrightarrow{\operatorname{Id}\otimes[\wedge]}\mathcal{E}_{X}\otimes\Omega_{X}^{[2]}$
vanishes. $\theta$ is called the Higgs field of $X$.
Lemma 2.4.
Let $\mathcal{E}_{X}$ be a locally free sheaf on a normal space $X$. Suppose that $\theta_{W}$ is a Higgs field of $\mathcal{E}_{X}|_{W}$ on an analytic subspace of codimension at least $2$, then $\theta_{W}$ extends to a Higgs field $\theta_{X}$ of $\mathcal{E}_{X}$.
Proof.
This is because $\operatorname{Hom}(\mathcal{E}_{X},\mathcal{E}_{X}\otimes\Omega_{X}^{[1]})$ and the target of the composed morphism $\theta\wedge\theta$ are reflexive. ∎
Lemma 2.5 (cf. [37, Lemma 4.12]).
Let $(\mathcal{E}_{X},\theta_{X})$ be a torsion-free Higgs sheaf on a normal space $X$. Suppose that $\mathcal{F}_{X}$ is a subsheaf of $\mathcal{E}_{X}$ and there exists a Zariski dense open subset $W\subset X_{\mathrm{reg}}$ such that $\mathcal{F}_{X}$ is $\theta_{X}$-invariant on $W$, then its saturation $\mathcal{F}^{sat}$ is generically $\theta_{X}$-invariant.
Remark 2.6.
Recall that the saturation of $\theta_{X_{\mathrm{reg}}}$ torsion-free subsheaves increases the slope (see e.g. [110, Lemma 2.30]) and are again $\theta_{X_{\mathrm{reg}}}$-invariant by Lemma 2.5. Thus, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-stable if and only if
(2.1) $\mu_{(\alpha_{0},\cdots,\alpha_{n-2})}(\mathcal{F}_{X})<\mu_{(\alpha_{0},\cdots,\alpha_{n-2})}(\mathcal{E}_{X}).$
holds for any saturated subsheaf $\mathcal{F}_{X}\subsetneq\mathcal{E}_{X}$ that is $\theta_{X_{\mathrm{reg}}}$-invariant. Ditto for semistability and polystability.
Combining Lemma 2.5 and Remark 2.6, it can be easily seen that
Lemma 2.7.
A torsion-free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ on a compact normal space $X$ is stable (resp. semistable) in the sense of [37, Definition 4.13] if and only if $(\mathcal{E}_{X},\theta_{X})|_{X_{\mathrm{reg}}}$ is stable (resp. semistable).
Remark 2.8.
Building on Lemma 2.4 and 2.7, all statements proved in the remainder of this subsection also hold for torsion-free Higgs sheaves over the whole space.
Following the spirits of [64, Pages 158], it can be easily seen that
Lemma 2.9.
Let $X$ be a compact complex normal space of dimension $n$, $\alpha_{0},\cdots,\alpha_{n-2}$ be nef classes and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ be a torsion-free Higgs sheaf on $X_{\mathrm{reg}}$. Suppose that
$0\rightarrow\mathcal{F}_{X}\rightarrow\mathcal{E}_{X}\rightarrow\mathcal{Q}_{X}\rightarrow 0$
is a short exact sequence of torsion-free sheaves such that $\mathcal{F}_{X_{\mathrm{reg}}}$ is $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$-invariant. If the induced Higgs sheaves $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}}),(\mathcal{Q}_{X_{\mathrm{reg}}},\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}})$ are $(\alpha_{0},\cdots,\alpha_{n-2})$-semistable and $\mu_{(\alpha_{0},\cdots,\alpha_{n-2})}(\mathcal{F}_{X})=\mu_{(\alpha_{0},\cdots,\alpha_{n-2})}(\mathcal{E}_{X})$, then $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-semistable.
2.2.2. Pull-back of Higgs sheaves
The existence of the pull-back of Higgs sheaves is essential and relies on Kebekus-Schnell’s work [63] of constructing pull-back of reflexive differentials.
Lemma 2.10 (Reflexive pull-back of Higgs field, cf. [110, Lemma 4.9]).
Let $f:Y\rightarrow X$ be a holomorphism between complex spaces with klt singularities and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ be a reflexive Higgs sheaf on $X_{\mathrm{reg}}$. Let $\mathcal{E}_{Y_{\mathrm{reg}}}:=(f^{*}\mathcal{E}_{X})^{\vee\vee}|_{X_{\mathrm{reg}}}$ and $\sigma$ be the non-locally-fre locus of $\mathcal{E}_{X_{\mathrm{reg}}}$. If $f^{-1}(X_{\mathrm{reg}})\cap Y_{\mathrm{reg}}\neq\emptyset$, then $\mathcal{E}_{Y_{\mathrm{reg}}}$ admits a Higgs field $\theta_{Y_{\mathrm{reg}}}$ such that $\theta_{Y_{\mathrm{reg}}}=f^{*}\theta_{X_{\mathrm{reg}}}$ on $f^{-1}(X_{\mathrm{reg}}\setminus\Sigma)$.
Notation 2.11 (Reflexive pull-back of Higgs sheaves).
We call $(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})$ the reflexive pull-back of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$, and denote it by $f^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$.
Remark 2.12.
If $\mathcal{E}_{X}$ is locally free, then $\mathcal{E}_{Y}=f^{*}\mathcal{E}_{X}$ is locally free and Lemma 2.4 implies $\theta_{Y_{\mathrm{reg}}}$ extends to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}$. We call $(\mathcal{E}_{Y},\theta_{Y})$ the pull-back of $(\mathcal{E}_{X},\theta_{X})$, and denote it by $f^{*}(\mathcal{E}_{X},\theta_{X})$.
2.3. Orbifold Chern classes
We refer to [90, Section 2] for a concise introduction of the definition of orbifold Chern classes via partial modifications [67]. A compact complex klt space has only quotient singularities in codimension 2 (cf. [31, Lemma 5.8]) and so admits a partial orbifold modification by [67], which allows us to define the orbifold first and second Chern classes as below.
Definition 2.13 (cf. [90, Section 2]).
Let $X$ be a compact complex klt space, $f:Y\rightarrow X$ be a partial orbifold resolution and $Y_{\mathrm{orb}}=\{V_{i},G_{i},\pi_{i}\}$ be the standard orbifold structure of $Y$. For any reflexive sheaf $\mathcal{E}$, $\widehat{c}_{2}(\mathcal{E})\in H_{2n-4}(X,\mathbb{R})$ is defined by
$\widehat{c}_{2}(\mathcal{E})\cdot\sigma:=c_{2}^{\mathrm{orb}}(f^{[*]}_{\mathrm{orb}}\mathcal{E})\cdot f^{*}\sigma,\ \forall\sigma\in H^{2n-4}(X,\mathbb{R})$
where $f^{[*]}_{\mathrm{orb}}\mathcal{E}=\{(f\circ\pi_{i})^{[*]}\mathcal{E}\}$ is a reflexive orbi-sheaf on $X_{\mathrm{orb}}$. Analogously, $\widehat{c}_{1}(\mathcal{E})\in{H_{2n-2}(X,\mathbb{R})}$ and $\widehat{c}_{1}^{2}(\mathcal{E})\in H_{2n-4}(X,\mathbb{R})$ can be introduced.
Definition 2.13 is independent of the choice of $f$. we refer to [90, Section 2] for a detailed discussion. When $\mathcal{E}=\mathcal{T}_{X}$, it coincides with the definition of $\widehat{c}_{2}(X)$ in [75, 31, 14].
Remark 2.14.
Here, $c_{2}^{\mathrm{orb}}(\cdot)$ denotes the orbifold Chern character in Bott–Chern cohomology for general coherent orbifold sheaves, introduced in [78]; for vector bundles this coincides with the definition via metrics given in [77]. The Grothendieck–Riemann–Roch formula for orbifold embeddings is also established in [78].
The following proposition is analogous to [55, Lemma 2.13] for $\mathbb{Q}$-Chern classes.
Proposition 2.15.
Let $\omega_{1},\cdots,\omega_{n-2}$ be n2n-2 nef and big classes on a compact complex klt space $X$ of dimension $n$. Then for any short exact sequence of coherent sheaves
$0\rightarrow\mathcal{F}\rightarrow\mathcal{E}\rightarrow\mathcal{Q}\rightarrow 0$
on $X$ with $\mathcal{E}$ reflexive and $\mathcal{Q}$ torsion-free, we have
(2.2) $\widehat{\mathrm{ch}}_{2}(\mathcal{E})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\leq\widehat{\mathrm{ch}}_{2}(\mathcal{F})\cdot[\omega_{1}]\cdots[\omega_{n-2}]+\widehat{\mathrm{ch}}_{2}(\mathcal{Q}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}],$
and if $\omega_{0},\cdots,\omega_{n-2}$ are Kähler and the equality in (2.2) holds then the natural map $\mathcal{Q}\rightarrow\mathcal{Q}^{\vee\vee}$ is an isomorphism in codimension $2$.
Proof of Proposition 2.15 essentially follows by considering the equality case of the argument of [53, Proposition 4.23] concerning $\widehat{c}_{2}(\mathcal{E})$ (cf. [62, Lemma 3.2] and [55, Lemma 2.13] for the projective case). We recall the following necessary lemmas in our proof.
Lemma 2.16 (cf. [90, Lemma 2.2]).
Let $f:Y\rightarrow X$ be a projective bimeromorphism between compact complex spaces of dimension $n$. Suppose that $\sigma_{1},\sigma_{2}\in H_{2n-2k}(X,\mathbb{R})$ agrees outside $f^{-1}(Z)$ for some analytic subspace $Z\subset X$ of codimension at least $k$. Then $f_{*}\sigma_{1}=f_{*}\sigma_{2}\in H_{2n-2k}(Y,\mathbb{R})$.
Lemma 2.17 (cf. [110, Lemma 2.17]).
Let $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow X_{\mathrm{orb}}$ be a bimeromorphism between compact complex orbifolds. Suppose that $\mathcal{F}_{\mathrm{orb}},\mathcal{E}_{\mathrm{orb}}$ are coherent orbi-sheaves on $Y_{\mathrm{orb}},X_{\mathrm{orb}}$, respectively, such that $(g_{\mathrm{orb}})_{*}\mathcal{F}_{\mathrm{orb}}=\mathcal{E}_{\mathrm{orb}}$ in codimension $k$ and the indeterminacy locus of $g_{\mathrm{orb}}^{-1}$ has codimension at least $k+1$. Then $g_{*}\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{F}_{\mathrm{orb}})=\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}}),$ i.e. $\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{F}_{\mathrm{orb}})\cdot g^{*}\gamma=\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}})\cdot\gamma$ for any $\gamma\in H^{2n-2k}(X,\mathbb{R})$,
Lemma 2.18.
Applying Lemma 2.2 and 2.17 and adapt the notations of Lemma 2.2 and Definition 2.13, it holds that
$\forall\sigma\in H^{2n-4}(X,\mathbb{R}),\ \widehat{c}_{2}(\mathcal{E})\cdot\sigma:=c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot g^{*}f^{*}\sigma.$
Ditto for $\widehat{c}_{1}(\mathcal{E})$ and $\widehat{c}_{1}^{2}(\mathcal{E})$.
Proposition 2.19.
Suppose that $X$ is a compact complex normal space of dimension $n$ equipped with a Kähler form $\omega_{X}$ and $f:Y\rightarrow X$ is a bimeromorphism from a compact complex space $Y$ with only quotient singularities to $X$. Let $Z$ be the indeterminacy locus of $f^{-1}$ and $Y_{\mathrm{orb}}:=\{(V_{i},G_{i},\mu_{i})\}$ be a complex orbifold structure. Then for any $V\subset\subset X\setminus Z$, there exists an smooth orbifold semi-positive $(1,1)$-form $\omega_{V,\mathrm{orb}}\in[f^{*}_{\mathrm{orb}}\omega_{X}]\in H^{1,1}_{BC}(X_{\mathrm{orb}},\mathbb{R})$ such that $\omega_{V,\mathrm{orb}}$ is strictly positive on $f_{\mathrm{orb}}^{*}V:=\{(f\circ\pi_{i})^{*}U\}$.
This lemma follows from [7, Section 6.1]. Note that let $f$ be the identity map $Y\rightarrow Y$, we conclude that $Y_{\mathrm{orb}}$ is Kähler if $Y$ is Kähler.
Proof of Proposition 2.15.
Let $g:W\rightarrow X$ be a partial orbifold resolution by Lemma 2.2 and $W_{\mathrm{orb}}:=\{(V_{i},G_{i},\pi_{i})\}$ be the standard orbifold structure of $W$. Denote the indeterminacy locus of $g^{-1}$ by $Z$. Let $g_{i}:=g\circ\pi_{i}$, $\mathcal{E}_{\mathrm{orb}}:=g_{\mathrm{orb}}^{[*]}\mathcal{E}$ and $\mathcal{F}_{\mathrm{orb}}$ be the saturation of $\{\operatorname{image}(g_{i}^{[*]}\mathcal{F}\rightarrow g_{i}^{[*]}\mathcal{E})\}$, where $g_{i}^{[*]}\mathcal{F}\rightarrow g_{i}^{[*]}\mathcal{E}$ is induced by the inclusion morphism $\mathcal{F}\rightarrow\mathcal{E}$. Since $\pi_{i}$ is étale in codimension $1$, $\mathcal{F}_{\mathrm{orb}}=g_{\mathrm{orb}}^{[*]}\mathcal{F}$ outside $g_{\mathrm{orb}}^{-1}(Z)$ due to the reflexivity. Then we have
(2.3) $\begin{split}(\widehat{\mathrm{ch}}_{2}(\mathcal{E})-\widehat{\mathrm{ch}}_{2}(\mathcal{F}))\cdot[\omega_{1}]\cdots[\omega_{n-2}]=&(\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}})-\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{F}_{\mathrm{orb}}))\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]\\ =&\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}}/\mathcal{F}_{\mathrm{orb}})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}],\end{split}$
where the first equality follows from the definition 2.13 and Lemma 2.17.
Let $\mathcal{Q}_{\mathrm{orb}}:=\mathcal{E}_{\mathrm{orb}}/\mathcal{F}_{\mathrm{orb}}$. $\mathcal{T}_{\mathrm{orb}}:=(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}/\mathcal{Q}_{\mathrm{orb}})$ is a torsion orbi-sheaf with the support $S_{\mathrm{orb}}$. For any open subset $V$ compactly supported in $Y\setminus g^{-1}(Z)$, Proposition 2.19 implies that there exists a smooth orbifold semi-positive $(1,1)$-form $\omega_{i,\mathrm{orb}}\in[g_{\mathrm{orb}}^{*}\omega_{i}]$ such that $\omega_{i,\mathrm{orb}}$ is strictly positive on $V_{\mathrm{orb}}$. Building on the GRR formula for orbifold embeddings (cf. [78, Theorem 1.1]) and the fact that $\mathrm{codim}(S_{\text{orb}})\geq 2$, it follows that
(2.4) $\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}/\mathcal{Q}_{\mathrm{orb}})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]\geq 0.$
Moreover, equality forces $V_{\text{orb}}\cap S_{\text{orb}}=\emptyset$, as can be shown by adapting the reasoning in [5, Section 7]. This is shown as follows. First, we take an orbifold embedded resolution $\pi_{\mathrm{orb}}:(\widehat{X}_{\mathrm{orb}},\widehat{\mathcal{S}}_{\mathrm{orb}})\rightarrow(X_{\text{orb}},Z_{\text{orb}})$ of $(X_{\text{orb}},Z_{\text{orb}})$ and all irreducible components $\widehat{\mathcal{S}}_{i,\text{orb}}$ of $\mathcal{S}_{\text{orb}}$ smooth with transversal intersection, which exists by the functorial embedded resolution [105, Theorem 2.0.2]. Then, an application of the GRR formula for orbifold embeddings gives
$\pi^{*}\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{T}_{\mathrm{orb}})\cdot\pi^{*}[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots\pi^{*}[g_{\mathrm{orb}}^{*}\omega_{n-2}]=\sum\limits\operatorname{rank}(\mathcal{T}_{\mathrm{orb}}|_{{\mathcal{S}}_{i,\mathrm{orb}}})\int_{\mathcal{S}_{i,\mathrm{orb}}}\pi_{\mathrm{orb}}^{*}\omega_{1,\mathrm{orb}}\wedge\cdots\wedge\omega_{n-2,\mathrm{orb}}$
since the image of the intersection of $\widehat{\mathcal{S}}_{\mathrm{orb}}$ and the exceptional divisor of $\pi_{\mathrm{orb}}$ has codimension at least $3$.
Combining (2.3), (2.4), we obtain
$(\widehat{\mathrm{ch}}_{2}(\mathcal{E})-\widehat{\mathrm{ch}}_{2}(\mathcal{F}))\cdot[\omega_{1}]\cdots[\omega_{n-2}]=$ $(-\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}/\mathcal{Q}_{\mathrm{orb}})+\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}))\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]$
$\leq$ $\mathrm{ch}_{2}^{\mathrm{orb}}(\mathcal{Q}_{\mathrm{orb}}^{\vee\vee})\cdot[g_{\mathrm{orb}}^{*}\omega_{1}]\cdots[g_{\mathrm{orb}}^{*}\omega_{n-2}]$
$=$ $\widehat{\mathrm{ch}}_{2}(\mathcal{Q}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]$
where the last equality follows from the fact that $\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}=g_{\mathrm{orb}}^{\vee\vee}\mathcal{Q}^{\vee\vee}$ outside $g_{\mathrm{orb}}^{-1}(Z)$. Then the inequality was proved. Moreover, equality forces $(Y\setminus f^{-1}(Z))_{\mathrm{orb}}\cap S_{\mathrm{orb}}=\emptyset$ since the choice of $V$ is arbitrary. This implies that $0\rightarrow\mathcal{F}_{\mathrm{orb}}\rightarrow\mathcal{E}_{\mathrm{orb}}\rightarrow\mathcal{Q}_{\mathrm{orb}}^{\vee\vee}\rightarrow 0$ is a short exact sequence of orbi-sheaves over $Y\setminus f^{-1}(Z)$, and hence it induces a short exact of reflexive sheaves
$0\rightarrow\mathcal{F}\rightarrow\mathcal{E}\rightarrow\mathcal{Q}^{\vee\vee}\rightarrow 0$
on $X\setminus Z$ by [33, Lemma A.3 and Lemma A.4]. In particular, $\mathcal{Q}^{\vee\vee}=\mathcal{Q}$ outside $Z$. The proof is complete. ∎
2.4. Quasi-étale morphisms
Definition 2.20.
Let $\gamma:Y\rightarrow X$ be a morphism between normal complex spaces.
  • (1)
    $\gamma$ is called Galois if there exists a finite group $G\subset\mathrm{Aut}(Y)$ such that $\gamma:Y\rightarrow X$ is isomorphic to the quotient map.
  • (2)
    $\gamma$ is called quasi-étale if $\dim X=\dim Y$ and étale in codimension $1$.
We have the following elementary property (cf. [31, Proposition 5.6] for $\mathcal{E}=\mathcal{T}_{X}$), which is analogous to [37, Section 3.8] for $\mathbb{Q}$-Chern classes.
Proposition 2.21 (Orbifold Chern classes under quasi-étale morphisms).
Let $X$ be a compact klt space of dimension $n$ and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ and $\gamma:X^{\prime}\rightarrow X$ is a quasi-étale Galois morphism between complex spaces induced by a finite group $G\subset\mathrm{Aut}(X^{\prime})$. Then $Y$ is also klt and the following equality holds for any reflexive sheaf $\mathcal{E}$ and any classes $\sigma\in H_{2n-2}(X,\mathbb{R})$
$\widehat{c}_{2}(\gamma^{[*]}\mathcal{E})\cdot\gamma^{*}\sigma=\deg\gamma\cdot\widehat{c}_{2}(\mathcal{E})\cdot\sigma$
Ditto for $\widehat{c}_{1}(\mathcal{E})$ and $\widehat{c}_{1}^{2}(\mathcal{E})$.
Proof.
The fact that $X^{\prime}$ is klt follows from the argument of [66, Proposition 5.20]. Let $f:Y\rightarrow X$, $f^{\prime}:Y^{\prime}\rightarrow X^{\prime}$ be partial orbifold resolutions of $X,X^{\prime}$, respectively and let $Y_{\mathrm{orb}}:=\{V_{\alpha}^{\prime},G_{\alpha}^{\prime},\mu_{\alpha}^{\prime}\},Y_{\mathrm{orb}}^{\prime}:=\{V_{\beta}^{\prime},G_{\beta}^{\prime},\mu_{\beta}^{\prime}\}$ be the maximal standard orbifold structures of $Y^{\prime},Y$, respectively. Denote the adjoint of the indeternimacy locus of $f^{-1}$ and the image of the indeterminacy locus of $(f^{\prime})^{-1}$ under $\gamma$ by $Z$, which has codimension at least $3$. Recall Lemma 2.18, we may assume that there exists some orbi-bundles $E_{\mathrm{orb}},E_{\mathrm{orb}}^{\prime}$ on $Y_{\mathrm{orb}},Y_{\mathrm{orb}}^{\prime}$ such that $E_{\mathrm{orb}}=f_{\mathrm{orb}}^{*}\mathcal{E},E_{\mathrm{orb}}^{\prime}=f_{\mathrm{orb}}^{\prime*}\mathcal{E}^{\prime}$ outside $f^{-1}(Z),(\gamma\circ f^{\prime})^{-1}(Z)$, respectively, and
$c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\sigma=\widehat{c}_{2}(\mathcal{E})\cdot\sigma,\ c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}}^{\prime})\cdot(f^{\prime})^{*}\gamma^{*}\sigma=\widehat{c}_{2}(\gamma^{[*]}\mathcal{E})\cdot\gamma^{*}\sigma.$
Thus it suffices to show that
$c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\sigma=\deg(\gamma)\cdot c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}}^{\prime})\cdot(f^{\prime})^{*}\gamma^{*}\sigma.$
Let $g_{\alpha,\beta}=W_{\alpha,\beta}\rightarrow V_{\alpha}\times_{X}V_{\beta}^{\prime}$ be a functorial resolution of singularities (cf. [22, Theorem 3.10]), then there is a induced action $G_{\alpha}\times G_{\beta}$ on $W_{\alpha,\beta}$ such that $g_{\alpha,\beta}$ is $(G_{\alpha}\times G_{\beta})$-invariant, which induces an effective complex orbifold $W_{\mathrm{orb}}:=\{W_{\alpha,\beta},G_{\alpha}\times G_{\beta}\}$. Let $W$ be the underlying space and we have the following commute diagram
[…]
Since $\gamma$ is quasi-étale, $\{V_{\beta}^{\prime},GG_{\beta}^{\prime},h\circ\mu_{\beta}^{\prime}\}$ can be viewed as a non-maximal standard orbifold structure of $Y\setminus\gamma^{-1}(Z)$. Applying Lemma 2.1, we conclude that the singularities of $Y^{\prime}\times_{X}Y$ lies in $pr_{2}^{-1}(\gamma^{-1}(Z))$, and thus Lemma 2.17 implies that
$c_{2}^{\mathrm{orb}}(g^{*}pr_{2}^{*}E_{\mathrm{orb}})\cdot g^{*}pr_{2}^{*}f^{*}\sigma=c_{2}^{\mathrm{orb}}(g^{*})\cdot c_{2}^{\mathrm{orb}}(g^{*}pr_{1}^{*}E_{\mathrm{orb}}^{\prime})\cdot g^{*}pr_{1}^{*}f^{*}\sigma$
because $g^{*}pr_{2}^{*}E_{\mathrm{orb}}$ and $g^{*}pr_{2}^{*}E_{\mathrm{orb}}^{\prime}$ coincides outside the preimage of $Z$. Note that $pr_{1}$ is generically biholomorphic and $pr_{2}$ is finite with degree $\deg(\gamma)$. The proof is complete. ∎
Remark 2.22 (Locally free sheaf).
Using a similar argument of the proof of Proposition 2.21, we have that for any locally free sheaf $\mathcal{E}$, $\widehat{c}_{1}(\mathcal{E}),\widehat{c}_{1}^{2}(\mathcal{E}),\widehat{c}_{2}(\mathcal{E})$ equal the usual product with the topological Chern classes of $\mathcal{E}$ in the sense of homology class.
Lemma 2.23 (Comparison).
Let $\mathcal{E}$ be a torsion-free sheaf on a compact normal space $X$. Then the following statements holds.
  • (1)
    For any $\sigma\in H^{2n-2}(X,\mathbb{R})$ and finite quasi-étale morphism $\gamma:Y\rightarrow X$ between compact normal spaces, we have
    $c_{1}(\gamma^{*}\mathcal{F})\cdot\gamma^{*}\sigma=\deg\gamma\cdot c_{1}(\mathcal{F})\cdot\gamma.$
  • (2)
    If $X$ is klt, $c_{1}(\mathcal{F})=\widehat{c}_{1}(\mathcal{F})$.
Proof.
The statement (1) can be implied by a similar argument of Lemma 2.21. The statement (2) follows from that for any partial orbifold resolution of singularities $f:Y\rightarrow X$, $f_{*}c_{1}(f^{[*]}\mathcal{F})=c_{1}(\mathcal{F})$ (see e.g. [110, Lemma 2.31]) and Lemma [110, Lemma 2.37]. ∎
Lemma 2.24 (Stability under quasi-étale morphisms).
Let $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ be a torsion-free Higgs sheaf on a compact complex normal space $X$ of dimension $n$ with rational singularities and $\alpha_{0},\cdots,\alpha_{n-2}$ be nef classes. Suppose that $\gamma$ is a quasi-étale morphism induced by some group $G$. Then
  • (1)
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-stable if and only if $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $G$-stable with respect to $\gamma^{*}\alpha_{0},\cdots,\gamma^{*}\alpha_{n-2}$. Ditto for polystability.
  • (2)
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-semistable if and only if $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\gamma^{*}\alpha_{0},\cdots,\gamma^{*}\alpha_{n-2})$-semistable.
Proof.
Based on the above lemmas, the statement (1) follows directly from the similar argument [110, Proposition 4.14]. Regarding (2), it suffices to notice that the uniqueness of the maximal destablizing subsheaf $\mathcal{F}_{Y_{\mathrm{reg}}}$ of $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ implies that $\mathcal{F}_{Y_{\mathrm{reg}}}$ is $G$-equivariant and any $G$-equivariant $\gamma^{[*]}\theta_{X_{\mathrm{reg}}}$-invariant subsheaf of $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ descends to a $\theta_{X_{\mathrm{reg}}}$ subsheaf of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$. ∎
3. Pluri-harmonic metrics of Higgs sheaves over the regular loci
In this section, we prove the existence of pluri-harmonic metrics on polystable reflexive Higgs sheaves with vanishing orbifold Chern number conditions (1.3).
We introduce some basic definitions to fix the conventions as follows. The Hitchin-Simpson connection ([97]) is defined by
$\bar{\partial}_{\theta}:=\bar{\partial}_{E}+\theta,\quad D_{H,\theta}^{1,0}:=\partial_{H}+\theta^{*H},\quad D_{H,\theta}=\bar{\partial}_{\theta}+D_{H,\theta}^{1,0},$
where $\partial_{H}$ is the $(1,0)$-part of the Chern connection $D_{H}$ of $(E,\bar{\partial}_{E},H)$ and $\theta^{*H}$ is the adjoint of $\theta$ with respect to $H$. The curvature of Hitchin-Simpson connection is
$F_{H,\theta}=F_{H}+[\theta,\theta^{*H}]+\partial_{H}\theta+\bar{\partial}_{E}\theta^{*H},$
where $F_{H}$ is the curvature of $D_{H}$.
Definition 3.1 (Pluri-harmonic metric on Higgs bundles).
Let $(E,\bar{\partial},\theta)$ be a Higgs bundle over a complex manifold $M$. We say that a Hermitian metric $H$ on $(E,\bar{\partial},\theta)$ is pluri-harmonic if $F_{H,\theta}=0$.
The main result of this section is the following.
Theorem 3.2.
Let $X$ be a compact Kähler klt space of dimension $n$. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is semistable with respect to some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$, then the following orbifold Bogomolov-Gieseker inequality holds.
(3.1) $\left(2\widehat{c}_{2}(\mathcal{E}_{X})-\frac{r-1}{r}\widehat{c}_{1}^{2}(\mathcal{E}_{X})\right)\cdot[\omega_{1}]\cdots[\omega_{n-2}]\geq 0.$
If $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{pHiggs}_{X_{\mathrm{reg}}}$, then $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ admits a pluri-harmonic metric $H_{X\setminus\Sigma}$ outside an analytic subspace $\Sigma\subsetneq X$ with $\mathrm{codim}_{X}\Sigma\geq 2$ and $X_{\text{Sing}}\subseteq\Sigma$ such that $\mathcal{E}_{X_{\mathrm{reg}}}|_{X\setminus\Sigma}$ is locally free.
Remark 3.3.
When $X$ is smooth and $\mathcal{E}_{X_{\mathrm{reg}}}$, Theorem 3.2 builds on the Donaldson-Uhlenbeck-Yau theorem [24, 104] and its generalization to Higgs bundles [97]; further developments appears in [2, 3, 71, 84, 85, 59, 73, 7, 10, 56, 57], etc. We refer to the introduction of [90] for an overview of developments on the orbifold Bogomolov-Gieseker inequality.
3.1. Strategy of Theorem 3.2
In this subsection, we outline the strategy of proving Theorem 3.2. The basic idea is to adapt the Higgs version [73] of Bando-Siu’s argument to an orbifold modification $h:W\rightarrow X$ as in [28] for stable reflexive sheaves. The main ingredients are:
  1. (i)
    The construction of an orbifold Higgs field of the pull-back of a reflexive Higgs sheaf via Kebekus-Schnell’s work [63] on pulling back reflexive differentials (cf. Proposition 3.4).
  2. (ii)
    The uniform Sobolev inequality for the degenerating family of orbifold Kähler metrics (cf. Proposition 3.9), which follows from the smooth approximation argument in [28, Section 3] and uniform geometric estimates in [44, 45, 46].
Once these ingredients are in place, the subsequent proof at this stage can be reduced to study the limiting behavior of the Hermitian-Yang-Mills flow over a non-compact balanced manifold $(X\setminus\Sigma,\omega)$ and the strategy in [73, Section 4] originated from [97] remains valid.
3.1.1. Pull-back of the Higgs field over orbifold modifications
We first review the construction in [110, Section 5.1].
Proposition 3.4.
Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a reflexive Higgs sheaf on the regular locus of a compact complex klt space $X$ and $\omega_{0},\cdots,\omega_{n-2}$ be Kähler forms, there exists a projective bimeromorphism $f:Y\rightarrow X$ from a compact Kähler space $Y$ with only quotient singularities to $X$ with the following data.
  • (a)
    An complex effective orbifold structure $Y_{\mathrm{orb}}:=\{U_{\mu},G_{\mu},\pi_{\mu}\}$ of $Y$.
  • (b)
    An analytic subspace $\Sigma\subset X$ of codimension at least $2$ such that $f_{\mu}:=f\circ\pi_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$ for each $i$, $X$ is smooth and $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free outside $\Sigma$.
  • (c)
    A Higgs orbi-bundle $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ over $Y_{\mathrm{orb}}$ with
    (3.2) $c_{1}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\gamma=\widehat{c}_{1}(\mathcal{E}_{X})\cdot\gamma=c_{1}(\mathcal{F})\cdot\gamma,\ \forall\gamma\in H_{2n-2}(X,\mathbb{R})$
    and
    (3.3) ${\mathrm{ch}}_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\beta=\widehat{\mathrm{ch}}_{2}(\mathcal{E}_{X})\cdot\beta,\ \forall\beta\in H_{2n-4}(X,\mathbb{R}).$
    such that
    (3.4) $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})=\{f_{\mu}^{*}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\}$
    outside $f_{\mathrm{orb}}^{-1}(\Sigma):=\{f_{\mu}^{-1}(\Sigma)\}$. The equality 3.3 also holds for $\widehat{c}_{1}^{2}(\mathcal{E}_{X})$.
Proof.
Let $h:W\rightarrow X$ be a partial orbifold resolution given by [67] and $W_{\mathrm{orb}}:=\{(V_{\mu},G_{\mu},\pi_{\mu})\}$ be the standard orbifold structure of $W$. Set $\mathcal{E}_{\mathrm{orb}}=h_{\mathrm{orb}}^{[*]}\mathcal{E}_{X}:=\{\big((h\circ\pi_{\mu})^{*}\mathcal{E}_{X}\big)^{\vee\vee}\}.$ According to Lemma 2.2, there exists a projective orbifold morphism $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow W_{\mathrm{orb}}$ from a compact complex effextive orbifold $Y_{\mathrm{orb}}:=\{U_{\mu},G_{\mu},\pi_{\mu}\}$ to $W_{\mathrm{orb}}$ such that $E_{\mathrm{orb}}:=g^{T}_{\mathrm{orb}}\mathcal{E}_{\mathrm{orb}}=\{g_{\mu}^{*}\mathcal{E}/(\mathrm{torsion})\}$ is a vector orbi-bundle. Let $g$ be the induced morphism of $g_{\mathrm{orb}}$ from the quotient space $Y$ of $Y_{\mathrm{orb}}$ to $W$, then $f:=h\circ g$ is a projective bimeromorphism.
Building on Lemma 2.10, we conclude that there exists a Higgs field $\theta_{\mathrm{orb}}=\{\theta_{i}\}$ of $E_{\mathrm{orb}}$ such that $\theta_{i}|_{f_{i}^{-1}(X_{\mathrm{reg}})}$ is the pull-back of $\theta$ via $f_{i}$. We summarize the dimensional properties by the constructions:
  • (i)
    There exists an analytic subspace $Z\subset X$ of codimension at least $3$ such that $g$ is biholomorphic outside $h^{-1}(Z)$.
  • (ii)
    $W_{\mathrm{orb}}$ is standard, i.e. $\pi_{i}$ is finite and quasi-étale for each $i$.
  • (iii)
    $g_{i}$ is biholomorphic outside the preimage of the non-locally free locus $S_{i}$ of $\mathcal{E}_{i}$, where $S_{i}$ has codimension at least $3$ due to the reflexivity of $\mathcal{E}_{i}$.
Combining Lemma 2.18 and 2.23, the proof is complete by choosing $\Sigma=Z\cup X_{sing}\cup_{i}h_{i}(S_{i})$. ∎
3.1.2. Bando-Siu’s methods
Suppose that $\omega_{0},\cdots,\omega_{n-2}$ are Kähler forms on $X$. Recall that $Y_{\mathrm{orb}}$ admits an orbifold Kähler form $\omega_{Y_{\mathrm{orb}}}=\{\omega_{\mu}\}$ by Proposition 2.19. Set
(3.5) $\omega_{\mathrm{orb},k,\epsilon}:=f_{\mathrm{orb}}^{\ast}\omega_{k}+\epsilon\omega_{Y_{\mathrm{orb}}}=\{f_{\mu}^{*}\omega_{k}+\epsilon\omega_{\mu}\}$
for all $0\leq k\leq n-2$, where $f_{\mathrm{orb}}^{*}\omega_{k}:=\{f_{\mu}^{*}\omega_{k}\}$. Recall that for each $\epsilon$, there exists a smooth orbifold Hermitian metric $\omega_{\mathrm{orb},\epsilon}$ (see e.g. [80, Pages 279]) such that
(3.6) $\omega_{\mathrm{orb},\epsilon}^{n-1}=\omega_{\mathrm{orb},0,\epsilon}\wedge\cdots\wedge\omega_{\mathrm{orb},n-2,\epsilon}.$
Note that $\omega_{\mathrm{orb},\epsilon}$ is an orbifold balanced metric (i.e. $\partial\omega_{\mathrm{orb},\epsilon^{n-1}}=0$) that is not Kähler in general. Given a smooth Hermitian metric $\hat{H}$ on the Higgs orbi-bundle $(E_{\mathrm{orb}},\bar{\partial}_{E_{\mathrm{orb}}},\theta_{\mathrm{orb}})$ , it is easy to see that there exists a uniform constant $\hat{C}_{0}$ such that
(3.7) $\int_{Y_{\mathrm{orb}}}(|\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{\hat{H}}|_{\hat{H}}+|\theta|_{\hat{H},\omega_{\mathrm{orb},\epsilon}}^{2})\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\leq\hat{C}_{0},$
for all $\epsilon$. Consider the Hermtian-Yang-Mills flow on the Higgs orbi-bundle $(E_{\mathrm{orb}},\bar{\partial}_{E_{\mathrm{orb}}},\theta_{\mathrm{orb}})$ with the fixed initial metric $\hat{H}_{\mathrm{orb}}:=\{\hat{H}_{\mu}\}$ and with respect to the Hermitian metric $\omega_{\epsilon}$,
[…]
where $\lambda_{\epsilon}=\frac{2\pi}{\operatorname{Vol}(Y_{\mathrm{orb}},\omega_{\mathrm{orb},\epsilon})}\mu_{\omega_{\mathrm{orb},\epsilon}}(E_{\mathrm{orb}})$. Following Simpson’s argument ([97]), we obtain the long time existence and uniqueness of the solution to the above flow (3.8). Actually, the needed estimates are valid on each orbifold chart $U_{i}$, whereby computations agree with those in the manifold case (see e.g. [27] for considerations on the existence of Hermitian-Einstein metrics of stable vector orbi-bundles, and [110, Section 3] for general stable Higgs orbi-bundles).
In Section 3.2, building on the uniform Sobolev inequality (Proposition 3.9) and following Bando-Siu’s argument [3], by choosing a subsequence, we shall prove that $H_{\mathrm{orb},\epsilon}(t)$ converges to smooth metrics $H_{\mathrm{orb}}(t)=\{H_{\mu}(t)\}$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$-topology as $\epsilon\rightarrow 0$ and thus $H_{\mathrm{orb}}(t)$ can be descend to a long time solution $H(t)$ of the Hermitian-Yang-Mills flow on $(X\setminus\Sigma)\times[0,+\infty)$, i.e. $H(t)$ satisfies:
[…]
because $\widehat{H}_{\mu}$ is $G_{\mu}$-invariant and $f_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$, where $\omega$ is a smooth Hermitian metric on $X$ such that
(3.10) $\omega^{n-1}=\omega_{0}\wedge\cdots\wedge\omega_{n-2}$
and $\lambda=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X}).$
In Section 3.3, we shall study the limiting behavior of the Hermitian-Yang-Mills flow (3.8), which enable us to obtain the following theorem.
Theorem 3.5.
In the setting in Proposition 3.4, when $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then by choosing a subsequence, we have $H(t)\rightarrow H_{\infty}$ in $C^{\infty}_{loc}(X\setminus\Sigma)$ and
$\sqrt{-1}\Lambda_{\omega}(F_{H_{\infty},\theta_{X_{\mathrm{reg}}}})=\lambda\cdot\operatorname{Id}_{\mathcal{E}_{\mathrm{reg}}}$
on $X\setminus\Sigma$. In the semistable case, we have
(3.11) $\sup_{x\in X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}(F_{H(t),\theta_{X_{\mathrm{reg}}}})-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}^{2}\rightarrow 0.$
To establish the Donaldson-Uhlenbeck-Yau theorem in our context, what remains open is the admissibility of $H_{\infty}$. The primary obstacle is proving a uniform estimate of $\theta_{X_{\mathrm{reg}}}$, a necessary step that is currently not attainable (cf. the smooth case in [73, Section 3]).
Question 3.6.
For $t_{0}>0$, whether $\|\theta_{X_{\mathrm{reg}}}\|_{H(t),\omega}(t>t_{0})$ are uniformly bounded along the Hermitian-Yang-Mill flow $H(t)$?
3.1.3. Completion of proof of Theorem 3.2
Building on Theorem 3.5, the proof of Theorem 3.2 can be complete. It suffices to consider the stable case. We borrow the ideas of [10, 9] using the Hodge index Theorem.
Let $X$ be a compact Kähler klt space of dimension $n$ and $\omega_{0},\cdots,\omega_{n-2}$ be n1n-1 Kähler forms. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ be a $(\omega_{0},\cdots,\omega_{n-2})$-stable reflexive Higgs sheaf on the regular locus of $X$ with vanishing orbifold Chern numbers (1.3). We adapts the notations in Section 3.1.1 and Section 3.1.1. It follows from (3.2) that
(3.12) $\lambda=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f^{*}_{\mathrm{orb}}\omega)^{n}}\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f^{*}_{\mathrm{orb}}\omega)^{n}}\frac{\widehat{c}_{1}(\mathcal{E}_{X})\cdot[\omega_{0}]\cdots[\omega_{n-2}]}{\operatorname{rank}(\mathcal{E}_{X})}=0.$
Since $\omega_{\epsilon}^{n-1}=\omega_{0,\epsilon}\wedge\cdots\wedge\omega_{n-2,\epsilon}$, we have
$0$ $=(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\omega_{\epsilon})\wedge\omega_{0,\epsilon}\wedge\cdots\wedge\omega_{n-2,\epsilon}$
$=(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\wedge\omega_{0,\epsilon}\wedge\cdots\wedge\omega_{n-2,\epsilon}$
where $\gamma_{\epsilon}=\frac{\omega_{\epsilon}^{n}}{\omega_{0,\epsilon}^{2}\wedge\cdots\wedge\omega_{n-2,\epsilon}}$ ($\epsilon>0$) is a smooth function on $Y_{\mathrm{orb}}$. Recall the following elementary statement ([102]).
Lemma 3.7.
Let $V$ be a complex linear space of dimension $n$. Let $\omega_{0},\cdots,\omega_{n-2}$ be n1n-1 positive $(1,1)$-forms. We define a Hermitian form on $\Lambda^{1,1}V$ by
$Q(\alpha,\sigma):=*(\alpha\wedge\bar{\sigma}\wedge\omega_{1}\wedge\cdots\wedge\omega_{n-2})$
and the mixed primitive subspace $P^{1,1}$ by
$P^{1,1}:=\{\alpha\in\Lambda^{1,1}V:\alpha\wedge\omega_{0}\wedge\cdots\wedge\omega_{n-2}=0\}.$
Then $Q$ is negative definite on $P^{1,1}$.
By applying Lemma 3.7 to $\sqrt{-1}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}(\sqrt{-1}F_{H_{\mathrm{orb},\epsilon}(t)})}{n}\gamma_{\epsilon}\omega_{0,\epsilon}$ with respect to $\omega_{0,\epsilon},\cdots,\omega_{n-2,\epsilon}$, we conclude that
$\operatorname{tr}\left((F_{H_{\mathrm{orb},\epsilon},\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\wedge(F_{H_{\mathrm{orb},\epsilon},\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\right)\wedge\Omega_{\epsilon}$
is positive outside $f_{\mathrm{orb}}^{-1}(\Sigma)$ for $0<\epsilon\ll 1$, and note that its descend via $f_{\mathrm{orb}}$ converges to
$\operatorname{tr}\left((F_{H(t),\theta_{X_{\mathrm{reg}}}}-\frac{\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}}{n}\cdot\gamma_{0}\omega_{0})\wedge(F_{H(t),\theta_{X_{\mathrm{reg}}}}-\frac{\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}}{n}\cdot\gamma_{0}\omega_{0})\right)\wedge\Omega$
as $\epsilon\rightarrow 0$ in $C^{\infty}_{loc}(X\setminus\Sigma)$, where $\Omega_{\epsilon}=\omega_{1,\epsilon}\wedge\cdots\wedge\omega_{n-2,\epsilon}$ and $\Omega=\omega_{1}\wedge\cdots\wedge\omega_{n-2}$. Combining (3.3), Remark 2.14, Lemma 3.15 and the Fatou Lemma, we have
$-8\pi^{2}\widehat{\mathrm{ch}}_{2}(\mathcal{E}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]=-8\pi^{2}{\mathrm{ch}}_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot[f_{\mathrm{orb}}^{*}\omega_{1}]\cdots[f_{\mathrm{orb}}^{*}\omega_{n-2}]$
$=$ $-8\pi^{2}\lim\limits_{\epsilon\rightarrow 0}\mathrm{ch}_{2}(E_{\mathrm{orb}})\cdot[\omega_{1,\epsilon}]\cdots[\omega_{n-2,\epsilon}]=\lim\limits_{\epsilon\rightarrow 0}\int_{Y_{\mathrm{orb}}}\operatorname{tr}(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}\wedge F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}})\wedge\Omega_{\epsilon}$
$=$ $\lim\limits_{\epsilon\rightarrow 0}\int_{Y_{\mathrm{orb}}}\operatorname{tr}\left((F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\wedge(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\right)\wedge\Omega_{\epsilon}$
$-\frac{1}{n^{2}}\int_{Y_{\mathrm{orb}}}|\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}|^{2}\gamma_{\epsilon}^{2}\omega_{0,\epsilon}^{2}\wedge\Omega_{\epsilon}$
$\geq$ $\int_{X\setminus\Sigma}\operatorname{tr}\left((F_{H(t),\theta_{X_{\mathrm{reg}}}}-\frac{\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}}{n}\cdot\gamma_{0}\omega_{0})\wedge(F_{H(t),\theta_{X_{\mathrm{reg}}}}-\frac{\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}}{n}\cdot\gamma_{0}\omega_{0})\right)\wedge\Omega$
$-\frac{1}{n^{2}}\int_{X\setminus\Sigma}|\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}|^{2}\gamma_{0}\omega^{n}.$
Similarly, let $t\rightarrow+\infty$, it follows from Theorem 3.5, (3.12) and (3.11) that
$0\geq\int_{X\setminus\Sigma}\operatorname{tr}(F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\wedge F_{H_{\infty},\theta_{X_{\mathrm{reg}}}})\wedge\omega_{1}\wedge\cdots\wedge\omega_{n-2}.$
Since $F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\wedge\omega_{1}\wedge\cdots\wedge\omega_{n-2}=\Lambda_{\omega}F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\cdot\omega^{n-1}=0$ outside $\Sigma$, it follows from Lemma 3.7 again that $F_{H_{\infty},\theta_{X_{\mathrm{reg}}}}\equiv 0$ outside $\Sigma$.
To complete the proof, it remains to prove Bogomolov-Gieseker inequality (3.1). Recall [110, Corollary 1.3] that for a $(\omega_{0},\cdots,\omega_{n-2})$-semistable reflexive Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})$,
$\mathcal{G}_{X_{\mathrm{reg}}}:=\operatorname{End}(\mathcal{E}_{X_{\mathrm{reg}}})=((\mathcal{E}_{X_{\mathrm{reg}}})^{\vee}\otimes\mathcal{E}_{X_{\mathrm{reg}}})^{\vee\vee}$
with the induced Higgs sheaf $\theta_{\mathcal{G}_{X_{\mathrm{reg}}}}$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable. It follows from [110, Lemma 2.45] that $\widehat{c}_{1}(\mathcal{G}_{X})=0$ and
(3.13) $\widehat{\Delta}(\mathcal{E}_{X})=\frac{1}{\operatorname{rank}(\mathcal{E}_{X})^{2}}\widehat{\Delta}(\mathcal{G}_{X})=-\frac{1}{\operatorname{rank}(\mathcal{E}_{X})}\widehat{\mathrm{ch}}_{2}(\mathcal{G}_{X})$
Then the Bogomolov-Gieseker inequality (3.1) can be concluded by applying the above argument to $(\mathcal{G}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$.
3.2. $C_{loc}^{\infty}$-convergence of $H_{\mathrm{orb},\epsilon}(t)\rightarrow H(t)$
Recall that $H_{\mathrm{orb},\epsilon}(t)$ is the long time solution of the Hermitian-Yang-Mills flow (3.8) on the Higgs orbi-bundle $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ with the fixed smooth initial metric $\hat{H}_{\mathrm{orb}}$ and with respect to the Hermitian metric $\omega_{\mathrm{orb},\epsilon}$. We shall prove the following statement.
Proposition 3.8.
By choosing subsequences, $H_{\mathrm{orb},\epsilon}(t)$ converges to $\{f_{\mu}^{*}H(t)\}$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$, where $H(t)$ is a smooth solution of the evolution equation (3.9) on $(X\setminus\Sigma)\times[0,+\infty)$ in $C_{loc}^{\infty}$-topology as $\epsilon\rightarrow 0$. Furthermore, $H(t)$ satisfies:
(3.14) $\int_{X\setminus\Sigma}|\Phi(H(t),\omega)|_{H(t)}\frac{\omega^{n}}{n!}\leq\int_{X\setminus\Sigma}|\Phi(\hat{H},\omega)|_{\hat{H}}\frac{\omega^{n}}{n!}\leq\hat{C}_{1},$
and
(3.15) $|\Phi(H(t+s),\omega)|_{H(t+t_{0})}(x)\leq\tilde{C}_{1}(\frac{1}{s^{2n+1}}+1)\int_{X\setminus\Sigma}|\Phi(H(t),\omega)|_{H(t)}\frac{\omega^{n}}{n!}$
for all $x\in X\setminus\Sigma$, $s>0$ and $t\geq 0$, where $\Phi(H,\omega)=\sqrt{-1}\Lambda_{\omega}(F_{H}+[\theta,\theta^{\ast H}])-\lambda_{\omega}\operatorname{Id}_{E}.$
3.2.1. Uniform Sobolev inequality for $\omega_{i,\epsilon}$
Thanks to Guo-Phong-Song-Sturm’s uniform geometric estimates ([44, 45, 46]) and the argument of [28, Section 3], we have the following uniform Sobolev inequalities for orbifold Kähler metrics $\omega_{\mathrm{orb},k,\epsilon}=f_{\mathrm{orb}}^{*}\omega_{k}+\epsilon\omega_{Y_{\mathrm{orb}}}$ (see (3.5)).
Proposition 3.9.
For any $q\in(1,\frac{n}{n-1})$, there exists a uniform constant $C_{S}>0$ independent of $\epsilon$ such that
(3.16) $(\int_{Y_{\mathrm{orb}}}|u|^{2q}\frac{\omega_{\mathrm{orb},k,\epsilon}^{n}}{n!})^{\frac{1}{q}}\leq C_{S}\left(\int_{Y_{\mathrm{orb}}}(|du|_{\omega_{\mathrm{orb},k,\epsilon}}^{2}+|u|^{2})\frac{\omega_{\mathrm{orb},k,\epsilon}^{n}}{n!}\right)$
for any $u\in L_{1}^{2}(Y_{\mathrm{orb}},\omega_{\mathrm{orb},k,\epsilon})$ and $0<\epsilon\leq 1$.
First, let us recall the uniform Sobolev inequality proved in [45, Section 4] (see also [91, Section 2.3] for general compact Hermitian manifolds).
Lemma 3.10.
[…]
$e^{F_{\omega}}\geq\gamma,\ \|e^{F_{\omega}}\|_{L^{1+\delta}}\leq K,\ [\omega]\cdot[\theta_{\widehat{Y}}]\leq A\text{ and }[\omega]^{n}\geq A^{-1}$
where $F_{\omega}:=\log(\frac{1}{V_{\omega}}\frac{\omega^{n}}{\theta_{\widehat{Y}}^{n}})$ for $V_{\omega}:=[\omega]^{n}$. For any $q\in(1,\frac{n}{n-1})$, there exists some constant $C=C(\gamma,A,K,\delta,q)>0$ such that for any $\omega\in\mathcal{K}(\widehat{Y},\theta_{\widehat{Y}},\gamma,A,K,\delta)$, we have that
$(\int_{\widehat{Y}}|u|^{2q}\omega^{n})^{\frac{1}{q}}\leq C_{S}\left(\int_{\widehat{Y}}(|du|_{\omega}^{2}+|u|^{2})\frac{\omega^{n}}{n!}\right)$
for any $u\in L_{1}^{2}(\widehat{Y},\omega)$.
Building on Lemma 3.10, Proposition 3.9 can be directly concluded by the argument using smooth approximation from [28, Section 3]. For reader’s convenience, we include a brief explanation.
Proof of Proposition 3.9.
For simplicity, we omit the subscript ‘k’ at the moment. Suppose that $\pi_{\mu}:U_{\mu}\rightarrow U_{\mu}/G_{\mu}$ is étale outside a proper analytic subspace $Z$. Recall Proposition 2.19 that there exists some smooth Kähler form $\omega_{0}$ on the underlying space $Y$ such that $\omega_{Y_{\mathrm{orb}}}:=\{\pi_{\mu}^{*}\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi_{\mu}\}$ for some $G_{\mu}$-invariant smooth function $\varphi_{\mu}$ on every chart $U_{\mu}$. Then $\varphi_{\mu}$ induces a bounded continous function $\varphi$ on $Y$ that is smooth outside $Z$ and in particular, $\omega_{Y_{\mathrm{orb}}}=\{\pi_{\mu}^{*}\omega_{Y}\}$ for $\omega_{Y}=\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi$. Thus, $\omega_{\mathrm{orb},\epsilon}:=\{f_{\mathrm{orb}}^{*}\omega+\epsilon(\pi_{\mu}\omega_{0}+\sqrt{-1}\partial\bar{\partial}\varphi_{\mu})\}$ can also be viewed as a Kähler current $\omega_{\epsilon}=\omega+\epsilon\omega_{0}+\epsilon\sqrt{-1}\partial\bar{\partial}\varphi$ on the underlying complex space $Y$ that is smooth outside $Z$. Let $f:\widehat{Y}\rightarrow Y$ be a resolution of singularities of $Y$ and $D$ be the exceptional divisor of $f\circ g$, we use the notation of Proposition 3.9 and Lemma 3.10.
As in [45, Section 7], the basic strategy is considering an approximation of $f^{*}\omega_{\epsilon}$ by a family $\{\omega_{\epsilon,j}\}_{j}$ of smooth Kähler metrics on $\widehat{Y}$. However, the situation here is slightly different to [45, Section 7] because $\log(\frac{f^{*}\omega_{\epsilon}^{n}}{\theta_{Y}^{n}})$ does not have log type analytic subvarieties (see [45, Definition 7.2]) in general, and is indeed the sum of a function $F_{\log}$ with log type singularities and a continuous function $G_{\epsilon}$ which is smooth outside $D$ (see [28, Lemma 3.9]). The main consideration in [28, Section 3] is addressing this issue. According to [28, Lemma 3.12, Lemma 3.13 and Lemma 3.15], there exists a sequence $\omega_{\epsilon,j}$ of Kähler metrics on $Y$ such that $\omega_{j,\epsilon}\in\mathcal{K}(Y,\theta_{Y},\gamma,A,K,\delta)$ for some $A,K,\delta$ and $\gamma$ independent of $j,\epsilon$, and for any integer $k\geq 0$, we have $\omega_{\epsilon,j}\rightarrow f^{*}\omega_{\epsilon}$ in $C_{loc}^{k}(Y\setminus D)$. Then by applying the same argument of [45, Pages 34] and Lemma 3.10, the uniform Sobolev inequality (3.16) for $\omega_{\epsilon}$ can be proved because $\mu_{\alpha}$ is unramified outside $Z$. ∎
It’s easy to see that there exists some constant $C_{k}>0$ independent of $\epsilon$ such that
(3.17) $C_{k}^{-1}\omega_{\mathrm{orb},k,\epsilon}\leq\omega_{\mathrm{orb},\epsilon}\leq C_{k}\omega_{\mathrm{orb},k,\epsilon}$
from the construction (see (3.5) and (3.6)). We have following uniform Sobolev inequalities for the Hermitian metrics $\omega_{\mathrm{orb},\epsilon}$ by applying Proposition 3.9 to $q=\frac{2n}{2n-1}$.
Proposition 3.11.
There exists a uniform constant $C_{S}>0$ such that
(3.18) $(\int_{Y_{\mathrm{orb}}}|u|^{\frac{4n}{2n-1}}\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!})^{\frac{2n-1}{2n}}\leq C_{S}\Big(\int_{Y_{\mathrm{orb}}}(|du|^{2}_{\omega_{\mathrm{orb},\epsilon}}+|u|^{2})\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\Big)$
for all $u\in L_{1}^{2}(Y_{\mathrm{orb}},\omega_{\mathrm{orb},\epsilon})$ and all $0<\epsilon\leq 1$.
3.2.2. Mean value inequality on orbifolds
We have the following standard lemma based on the Stokes formula for orbifolds (see e.g. [8, Theorem 3.4.2]) and the fact that $d\omega_{\mathrm{orb}}^{n-1}=0$.
Lemma 3.12.
Let $(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$ be an $n$-dimensional compact complex Gauduchon orbifold. Suppose that the nonnegative function $\phi(x,t)\in C^{\infty}(Y_{\mathrm{orb}}\times[0,\infty))$ is a solution of the following heat flow
(3.19) $(2\sqrt{-1}\Lambda_{\omega_{\mathrm{orb}}}\partial\overline{\partial}-\frac{\partial}{\partial t})\phi\geq 0.$
Then
$\|\phi(t)\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}\leq\|\phi(t_{0})\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}$
for any $t>t_{0}\geq 0$.
The following lemma can be concluded by running the argument of [12, Appendix] in the context of orbifolds. The only difference is that the constant $\widehat{C}_{1}$ doesn’t depend on $d\omega_{\mathrm{orb}}^{n-1}$ since we assume that $\omega_{\mathrm{orb}}$ is balanced (see [12, (A.27)].
Lemma 3.13.
Let $(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$ be an $n$-dimensional compact balanced orbifold. Suppose that $\phi(x,t)\in C^{\infty}(Y_{\mathrm{orb}}\times[0,\infty))$ is a nonnegative function satisfying the following inequality
(3.20) $(2\sqrt{-1}\Lambda_{\omega_{\mathrm{orb}}}\partial\overline{\partial}-\frac{\partial}{\partial t})\phi\geq-\breve{C}\phi$
with $\phi(0)\in L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$, $\phi(t)\in L^{\infty}(X\setminus\Sigma,\omega)$ and $\|\phi(t)\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}\leq\|\phi(t_{0})\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}$ for any $t>t_{0}\geq 0$. Then we have the following statement.
(1) For any $s>0$ and $t_{0}\geq 0$,
(3.21) $\|\phi(t_{0}+s)\|_{L^{\infty}(Y_{\mathrm{orb}})}\leq\hat{C}_{1}(\frac{1}{s^{2n+1}}+1)\|\phi(t_{0})\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})},$
where $\hat{C}_{1}$ is a positive constant depending only on $n$, $\breve{C}$ and the uniform Sobolev constant $C_{S}(\omega_{\mathrm{orb}})$.
(2) Let $Z_{\mathrm{orb}}\subset Y_{\mathrm{orb}}$ be a closed subset and $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus\tilde{\Sigma}_{\mathrm{orb}}$ and $\hat{\omega}_{\mathrm{orb}}$ be an another orbifold Hermitian metric on $Y_{\mathrm{orb}}$. If there exist positive constants $\delta_{0}$ and $C_{\bar{\Omega}_{|}orb}$ such that $\overline{B_{\delta_{0}}(\bar{\Omega})}\subset Y_{\mathrm{orb}}\setminus\Sigma_{\mathrm{orb}}$ and $C_{\bar{\Omega}_{\mathrm{orb}}}^{-1}\hat{\omega}_{\mathrm{orb}}\leq\omega_{\mathrm{orb}}\leq C_{\bar{\Omega}_{\mathrm{orb}}}\hat{\omega}_{\mathrm{orb}}$ on $\overline{B_{\delta_{0}}(\bar{\Omega}_{\mathrm{orb}})}$, where $\overline{B_{\delta}(\bar{\Omega}_{\mathrm{orb}})}$ is the $\delta$-neighborhood of $\bar{\Omega}_{\mathrm{orb}}$ with respect to $\hat{\omega}_{\mathrm{orb}}$. Then,
(3.22) $\sup_{\overline{B_{\delta_{1}}(\bar{\Omega}_{\mathrm{orb}})}\times[0,T]}\phi(x,t)\leq\hat{C}_{2}((\delta_{2}-\delta_{1})^{-1}+1)^{2n+1}(\int_{0}^{T}\|\phi(t)\|_{L^{1}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})}dt+\|\phi(0)\|_{L^{\infty}(\overline{B_{\delta_{2}}(\bar{\Omega}_{\mathrm{orb}})})})$
for all $0\leq\delta_{1}<\delta_{2}\leq\delta_{0}$, where $\hat{C}_{2}$ is a constant depending only on $n$, $\breve{C}$, the Sobolev constant $C_{S}(\omega_{\mathrm{orb}})$ and $\operatorname{Vol}(Y_{\mathrm{orb}},\omega_{\mathrm{orb}})$.
3.2.3. Proof of Proposition 3.8
The following estimates are proved by Simpson for manifolds and the proof is same by restriction to the orbifold chart $U_{i}$ ([97, Lemma 6.1]).
Lemma 3.14.
Along the heat flow (3.8), we have
(3.23) $(\Delta_{{\omega_{\mathrm{orb},\epsilon}}}-\frac{\partial}{\partial t})\operatorname{tr}(\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon}))=0,$
(3.24) $(\Delta_{\omega_{\mathrm{orb},\epsilon}}-\frac{\partial}{\partial t})|\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t)}^{2}=2|D_{H_{\mathrm{orb},\epsilon},\phi}(\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon}))|^{2}_{H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon}},$
and
(3.25) $(\Delta_{\omega_{\mathrm{orb},\epsilon}}-\frac{\partial}{\partial t})|\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t)}\geq 0.$
By (3.7) and (3.25), we have
(3.26) $\int_{Y_{\mathrm{orb}}}|\Phi(H_{\mathrm{orb},\epsilon}(t_{2}),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t_{2})}\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\leq\int_{Y_{\mathrm{orb}}}|\Phi(H_{\mathrm{orb},\epsilon}(t_{1}),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t_{1})}\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\leq\hat{C}_{1}$
for any $0\leq t_{1}\leq t_{2}$, where $\hat{C}_{1}$ is a constant independent of $\epsilon$. Since $\omega_{\mathrm{orb},\epsilon,k}$ is strictly positive outside $f_{\mathrm{orb}}^{-1}(\Sigma)$ and so does $\omega_{\mathrm{orb},\epsilon}$. Building on the uniform Sobolev inequalities (3.18), (3.17), (3.26) Lemma 3.12 and Lemma 3.13, we have
Lemma 3.15.
There exists constant $\tilde{C}_{1}$ which independent of $\epsilon$, such that
(3.27) $\||\Phi(H_{\mathrm{orb},\epsilon}(t_{0}+s),\omega_{\mathrm{orb},\epsilon})|_{H_{\epsilon}(t_{0}+s)}\|_{L^{\infty}(Y_{\mathrm{orb}})}\leq\tilde{C}_{1}(\frac{1}{s^{2n+1}}+1)\||\Phi(H_{\mathrm{orb},\epsilon}(t_{0}),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t_{0})}\|_{L^{1}(Y_{\mathrm{orb}})}$
for any $s>0$ and $t_{0}\geq 0$. Let $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)$, there exists a constant $\tilde{C}_{2}(\bar{\Omega}_{\mathrm{orb}})$ which independent of $\epsilon$, such that
(3.28) $|\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t)}(x)\leq\tilde{C}_{2}(\bar{\Omega}_{\mathrm{orb}})$
for any $(x,t)\in\bar{\Omega}_{\mathrm{orb}}\times[0,\infty)$.
Setting $\exp(s_{\mathrm{orb},\epsilon}(t))=h_{\mathrm{orb},\epsilon}(t)=\hat{H}_{\mathrm{orb}}^{-1}H_{\mathrm{orb},\epsilon}(t),$ we have
(3.29) $\begin{split}\frac{\partial}{\partial t}\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}^{-1}(t))=&\ \frac{\operatorname{tr}(h_{\mathrm{orb},\epsilon}(t)\cdot h_{\mathrm{orb},\epsilon}^{-1}(t)\frac{\partial h_{\mathrm{orb},\epsilon}(t)}{\partial t})-\operatorname{tr}(h_{\mathrm{orb},\epsilon}^{-1}(t)\frac{\partial h_{\mathrm{orb},\epsilon}(t)}{\partial t}\cdot h_{\mathrm{orb},\epsilon}^{-1}(t))}{\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}^{-1}(t)}\\ \leq&\ 2|\Phi(H_{\mathrm{orb},\epsilon}(t),\omega_{\mathrm{orb},\epsilon})|_{H_{\mathrm{orb},\epsilon}(t)},\end{split}$
and
(3.30) $\log(\frac{1}{2r}(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)^{-1}))\leq|s_{\epsilon}(t)|_{\hat{H}}\leq r^{\frac{1}{2}}\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)^{-1}),$
where $r=\operatorname{rank}(\mathcal{E}_{X})$. By (3.26) and (3.28), we have
(3.31) $\int_{Y_{\mathrm{orb}}}\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}^{-1}(t))-\log(2r)\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\leq 2\hat{C}_{1}t,$
(3.32) $\log(\operatorname{tr}h_{\mathrm{orb},\epsilon}(t)+\operatorname{tr}h_{\mathrm{orb},\epsilon}^{-1}(t))-\log(2r)\leq 2\tilde{C}_{2}(\bar{\Omega}_{\mathrm{orb}})t$
and
(3.33) $|s_{\mathrm{orb},\epsilon}(t)|_{\hat{H}}(x)\leq 2r^{\frac{1}{2}}\tilde{C}_{2}(\bar{\Omega}_{\mathrm{orb}})t+r^{\frac{1}{2}}\log(2r)$
for all $(x,t)\in\bar{\Omega}_{\mathrm{orb}}\times[0,\infty)$.
By using the uniform local $C^{0}$-estimate (3.33) and following the argument in [73, Lemma 2.5], we can derive uniform local $C^{\infty}$-estimates of $h_{\mathrm{orb},\epsilon}(t)$. The only difference is the Riemann curvature replacing by the Chern curvature of $\omega_{\mathrm{orb},\epsilon}$, so we omit the proof.
Lemma 3.16.
Let $\bar{\Omega}_{\mathrm{orb}}$ be a compact subset of $Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)$ and $T>0$. Then for every integer $j\geq 0$, there exists a constant $\bar{C}_{j+2}(\bar{\Omega}_{\mathrm{orb}},T)$ which independent of $\epsilon$, such that
(3.34) $\max_{(x,t)\in\bar{\Omega}_{\mathrm{orb}}\times[0,T]}|\nabla_{\hat{H}_{\mathrm{orb}}}^{j+2}h_{\mathrm{orb},\epsilon}(t)|_{\hat{H}_{\mathrm{orb}},\omega_{\mathrm{orb},\epsilon}}(x)\leq\hat{C}_{j+2}$
for all $0<\epsilon\leq 1$, where $\nabla_{\hat{H}_{\mathrm{orb}}}$ is the covariant derivative with respect to the Chern connection $D_{\hat{H}_{\mathrm{orb}}}$ of $\hat{H}$ and the Chern connection $D_{\omega_{\mathrm{orb},\epsilon}}$ of $\omega_{\mathrm{orb},\epsilon}$.
Note that $\omega_{\mathrm{orb},\epsilon}$ converges to $f_{\mathrm{orb}}^{*}\omega$ from the construction (3.5) in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$ and
$\lambda=\lim\limits_{\epsilon\rightarrow 0}\lambda_{\epsilon}=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(f^{*}\omega_{0},\cdots,f^{*}\omega_{n-2})}(E_{\mathrm{orb}})=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})$
where the last equality follows from (3.2) and $\hat{H}$ denotes the descend of $\hat{H}_{\mathrm{orb}}|_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}$. From the above local uniform $C^{\infty}$-bounds on $H_{\mathrm{orb},\epsilon}(t)$, we get that, by choosing a subsequence, $H_{\mathrm{orb},\epsilon}(t)$ converges to smooth metrics $H_{\mathrm{orb}}(t)$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}},f_{\mathrm{orb}}^{-1}(\Sigma))$ as $\epsilon\rightarrow 0$, which satisfies (3.9). Since $f_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$, by descending $H_{\mathrm{orb}}(t)|_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}$ to a smooth Hermitian metric $H(t)$ of $\mathcal{E}_{X_{\mathrm{reg}}}|_{X\setminus\Sigma}$, the proof of Proposition 3.8 is complete.
3.3. Limiting behavior of $H(t)$
Note that $(X\setminus\Sigma,\omega)$ is a non-compact balanced manifold since $\omega^{n-1}=\omega_{0}\wedge\cdots\wedge\omega_{n-2}$. Then $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus\Sigma}$ can be seen as a Higgs bundle on $(X\setminus\Sigma,\omega)$. For simplicity, we simply write $\theta_{X_{\mathrm{reg}}}$ as $\theta$ at the moment. Let us recall some basic lemmas.
Lemma 3.17 (Cutoff functions, see e.g. [91, Lemma 5.6]).
Let $(X,\omega)$ be a compact Hermitian space of dimension $n$ and $\Sigma\subset X$ be an analytic subspace of codimension $\geq k$. Then there exists cutoff functions $\{\rho_{i}\}_{i\in N}$ ($supp(\rho_{i})\subset X\setminus\Sigma$) increasing to $1_{X\setminus\Sigma}$ and a positive constant $C$ such that $\int_{X}|d\rho_{i}|^{2k}_{\omega}+|\partial\bar{\partial}\rho_{i}|^{k}_{\omega}\omega^{n}\leq C$ for all $i\in N$ and
(3.35) $\lim_{i\rightarrow+\infty}\int_{X}|d\rho_{i}|^{2k-\delta}_{\omega}+|\partial\bar{\partial}\rho_{i}|^{k-\delta}_{\omega}\omega^{n}=0$
for any $\delta>0$ small.
Lemma 3.18.
In lightning of Proposition 3.18 and Lemma 3.17, the uniform Sobolev inequality holds for $(X\setminus\Sigma,\omega)$.
Using the above cutoff functions and Lemma 3.18, we obtain the following mean value inequality by applying the standard Moser’s iteration as in [73, Lemma 5.2] and the stokes formula.
Lemma 3.19 (Mean value formula).
Let $(X,\omega)$ be a compact balanced space of dimension $n$ and $\Sigma\subset X$ be an analytic subspace of codimension $\geq 2$. Suppose that $u\in C^{2}(X\setminus\Sigma)$ is a bounded positive function satisfying $\Delta_{\omega}u\geq-A_{0}$. Then there exists a positive constant $\widetilde{C}$ depends only on $C_{S}$, $A_{0}$ and $\operatorname{Vol}(X,\omega)=\int_{X_{\mathrm{reg}}}\frac{\omega^{n}}{n!}$ such that
$\|u\|_{L^{\infty}(X\setminus\Sigma)}\leq\widetilde{C}(\|u\|_{L^{1}(X\setminus\Sigma,\omega)}+1).$
Lemma 3.20 (Stokes formula).
Let $(X,\omega)$ be a compact Hermitian space of dimension $n$ and $\Sigma\subset X$ be an analytic subspace of codimension $\geq 2$. Suppose $\eta$ is a (2n1)(2n-1)-form with $\int_{X\setminus\Sigma}|\eta|_{\omega}^{2}\frac{\omega^{n}}{n!}<\infty$. Then if $\mathrm{d}\eta$ is integrable, we have $\int_{X}\mathrm{d}\eta=0.$
These lemmas are used to take the roles of three assumptions in [97] and then we can follow a strategy of [97] as in [73, Section 4]. The remaining of the proof is a slight modification of [73, Section 4] to our context. We outline the argument for the completeness.
3.3.1. The Donaldson’s functional
Let $K$ and $H$ are two Hermitian metrics on $\mathcal{E}|_{X\setminus\Sigma}$, and set $\exp{s}=h=K^{-1}H$. By Lemma 3.1 in [97], we have
(3.36) $\Delta_{\omega}\log(\operatorname{tr}h+\operatorname{tr}h^{-1})\geq-2|\Lambda_{\omega}(F_{H,\theta})|_{H}-2|\Lambda_{\omega}(F_{K,\theta})|_{K},$
and
(3.37) $\Delta_{\omega,K}^{\prime}h=h\sqrt{-1}\Lambda_{\omega}(F_{K,\theta}-F_{H,\theta})-\sqrt{-1}\Lambda_{\omega}(\bar{\partial}_{\theta}h\wedge h^{-1}D_{K,\theta}^{1,0}h),$
where $\Delta_{\omega,K}^{\prime}=-\sqrt{-1}\Lambda_{\omega}\bar{\partial}_{\theta}D_{K,\theta}^{1,0}$.
Let $S_{K}=S_{K}(\mathcal{E}_{X}|_{X\setminus\Sigma})$ be the real vector bundle of $K$-self-adjoint endomorphisms of $\mathcal{E}|_{X\setminus\Sigma}$, and $\mathcal{P}(S_{K})$ be the normed space of smooth sections $s\in\Gamma(S_{K})$ with norm
(3.38) $\|s\|_{P}=\sup_{X\setminus\Sigma}|s|+\|\bar{\partial}_{\theta}s\|_{L^{2}}+\|\Delta_{\omega,K}^{\prime}s\|_{L^{1}}.$
Let $\mathcal{P}$ denote the space of smooth Hermitian metrics $K$ such that $\int_{X\setminus\Sigma}|\Lambda_{\omega}F_{K,\theta}|_{\omega}\frac{\omega^{n}}{n!}<\infty.$ Due to Proposition 4.1 in [97], we can make $\mathcal{P}$ into analytic manifold with local charts $s\mapsto Ke^{s}$, $s\in\mathcal{P}(S_{K})$. Let $\mathcal{P}_{0}$ be one of the components covered by these charts. Let’s recall the Donaldson’s functional defined on the space $\mathcal{P}_{0}$ (see Section 5 in [97] for details),
(3.39) $\mu_{\omega}(K,H)=\int_{X\setminus\Sigma}\operatorname{tr}(s\sqrt{-1}\Lambda_{\omega}F_{K,\theta})+\langle\Psi(s)(\overline{\partial}_{\theta}s),\overline{\partial}_{\theta}s\rangle_{K}\frac{\omega^{n}}{n!},$
where $\Psi(x,y)=(x-y)^{-2}(e^{y-x}-(y-x)-1)$.
The following proposition is standard because the argument of [97, Section 5] can be directly applied since $d\omega^{n-1}=0$.
Proposition 3.21.
Let $H(t)=Kh(t)$, $0\leq t\leq 1$, be a piecewise differentiable curve connecting metrics $K$ and $H$ in the space $\mathcal{P}_{0}$. If $d\omega^{n-1}\equiv 0$, then
(3.40) $\mu_{\omega}(K,H)=\int_{0}^{1}\int_{M\setminus\Sigma}\sqrt{-1}\operatorname{tr}(h^{-1}\frac{\partial h}{\partial t}\cdot F_{H(t),\theta})\wedge\frac{\omega^{n-1}}{(n-1)!}dt,$
and
(3.41) $\frac{d}{dt}\mu_{\omega}(K,H(t))=\int_{M\setminus\Sigma}\sqrt{-1}\operatorname{tr}(h^{-1}\frac{\partial h}{\partial t}\cdot F_{H(t),\theta})\wedge\frac{\omega^{n-1}}{(n-1)!}$
3.3.2. Approximate Hermitian-Einstein structure
Let $H(t)$ be the long time solution of (3.9) on the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})|_{X\setminus\Sigma}$. Since we have known that $|\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}|_{H(t)}$ is uniformly bounded for $t\geq t_{0}>0$, by (3.37) and Lemma 3.17, it is easy to see that $H(t)$ (for every $t>0$) belongs to the space $\mathcal{P}_{0}$. By (3.41), we have a formula for the derivative of Donaldson’s functional with respect to $t$,
(3.42) $\frac{d}{dt}\mu_{\omega}(H(t_{0}),H(t))=-2\int_{X\setminus\Sigma}|\Phi(H(t),\omega)|_{H(t)}^{2}\frac{\omega^{n}}{n!}.$
Set $\exp{s(t)}=h(t)=\hat{H}^{-1}H(t)$ and $\exp{s(t_{1},t_{2})}=h(t_{1},t_{2})=H^{-1}(t_{1})H(t_{2}).$ By the inequalities (3.36), (3.30), (3.14), (3.15) and Lemma 3.19, we have
(3.43) $r^{-\frac{1}{2}}\|s(t_{1},t_{2})\|_{L^{1}(X\setminus\Sigma,\omega,H(t_{1}))}\leq 2\hat{C}_{1}(t_{2}-t_{1})+\operatorname{Vol}(M,\omega)\log(2r),$
(3.44) $\|s(t_{1},t_{2})\|_{L^{\infty}(X\setminus\Sigma,H(t_{1}))}\leq r^{\frac{1}{2}}\{2C_{S}\hat{C}_{1}(t_{0}^{-n}+1)(t_{2}-t_{1})+\log 2r\}$
and
(3.45) $\|s(t_{1},t_{2})\|_{L^{\infty}(X\setminus\Sigma,H(t_{1}))}\leq C_{1}(t_{0}^{-1})\|s(t_{1},t_{2})\|_{L^{1}(M\setminus\Sigma,\omega,H(t_{1}))}+C_{2}(t_{0}^{-1})$
for $0<t_{0}\leq t_{1}\leq t_{2}$, where $C_{1}(t_{0}^{-1})$ and $C_{2}(t_{0}^{-1})$ are constants depending only on $r$, the Sobolev constant $C_{S}$, $\hat{C}_{1}$, $t_{0}^{-1}$ and $\operatorname{Vol}(X,\omega)$.
Proposition 3.22.
Let $H(t)$ be the long time solution of (3.9) on the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})|_{X\setminus\Sigma}$ with initial metric $\hat{H}$, and suppose that
(3.46) $\int_{X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta_{X_{\mathrm{reg}}}}-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}\frac{\omega^{n}}{n!}\leq\hat{C}_{1},$
where $\hat{C}_{1}$ is a positive constant. Setting
(3.47) $\int_{X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}^{2}\frac{\omega^{n}}{n!}\rightarrow C^{\ast}\geq 0,$
as $t\rightarrow+\infty$. Assume that there is a sequence $t_{i}$ such that
(3.48) $\|s(1,t_{i})\|_{L^{1}(X\setminus\Sigma,H(1))}\rightarrow+\infty.$
Then the following statement holds.
  1. (i)
    There is a section $)u_{\infty}\in L_{1}^{2}(\Gamma(S_{\hat{H}}$)) such that $\operatorname{tr}u_{\infty}=0$, $\|u_{\infty}\|_{L^{2}}=1$ and the eigenvalues of $u_{\infty}$ are almost everywhere constants and not all equal. Furthermore, assume $\mu_{1}<\mu_{2}<\cdots<\mu_{l}$ ($l\geq 2$) are the distinct eigenvalues of $u_{\infty}$. Defining smooth functions $P_{\alpha}:\mathbb{R}\to\mathbb{R}$ such that
    (3.49) $P_{\alpha}(x)=\left\{\begin{split}\ 1,\quad&x\leq\mu_{\alpha},\\ \ 0,\quad&x\geq\mu_{{\alpha}+1}.\end{split}\right$.
    and setting $\pi_{\alpha}=P_{\alpha}(u_{\infty})$ for every $1\leq\alpha\leq l-1$. Then every $\pi_{\alpha}$ defines a saturated $\theta_{X_{\mathrm{reg}}}$-invariant subsheaf $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ of the Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$, and every $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ can be extended to a saturated subsheaf $\mathcal{E}_{\alpha}$ of $\mathcal{E}_{X}$.
  2. (ii)
    $2\pi\sum_{\alpha=1}^{l-1}(\mu_{\alpha+1}-\mu_{\alpha})\operatorname{rank}(\mathcal{E}_{\alpha})(\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})-\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{\alpha}))\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.$
Proof.
By (3.45) and the assumption (3.48), it is easy to check that
(3.50) $\lim_{i\rightarrow\infty}\|s(t_{0},t_{i})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}\rightarrow+\infty,$
and
[…]
for all $0<t_{0}\leq 1\leq t$, where $C_{3}$ and $C_{4}$ are constants depending only on the Sobolev constant $C_{S}$, $r$, $\hat{C}_{1}$ and the geometry of $(M\setminus\Sigma,\omega)$.
Setting $u_{i}(t_{0})=\|s(t_{0},t_{i})\|_{L^{1}}^{-1}s(t_{0},t_{i})\in S_{H(t_{0})}$, then $\|u_{i}(t_{0})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}=1$ and
(3.52) $\int_{X\setminus\Sigma}\operatorname{tr}u_{i}(t_{0})\frac{\omega^{n}}{n!}=\|s(t_{0},t_{i})\|_{L^{1}}^{-1}\int_{t_{0}}^{t_{i}}\int_{X\setminus\Sigma}\operatorname{tr}(\Phi(H(t),\omega))\frac{\omega^{n}}{n!}dt=0.$
On the other hand, By (3.42), we have
(3.53) $\mu_{\omega}(H(t_{0}),H(t))=-2\int_{t_{0}}^{t}\int_{X\setminus\Sigma}|\Phi(H(s),\omega)|_{H(s)}^{2}\frac{\omega^{n}}{n!}ds\leq-2C^{\ast}(t-t_{0})$
for all $0<t_{0}\leq t$. Then it is clear that (3.43) implies
(3.54) $\liminf_{t\rightarrow+\infty}\frac{-\mu_{\omega}(H(t_{0}),H(t))}{\|s(t_{0},t)\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}}\geq r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.$
By the inequalities (3.54), (3.48), (3.51) and the Lemma 5.4 in [97], we can see that, by choosing a subsequence which we also denote by $u_{i}(t_{0})$, we have $u_{i}(t_{0})\rightarrow u_{\infty}(t_{0})$ weakly in $L_{1}^{2}$ as $i\rightarrow\infty$, where the limit $u_{\infty}(t_{0})$ satisfies: $\|u_{\infty}(t_{0})\|_{L^{1}(M\setminus\Sigma,\omega,H(t_{0}))}=1$, $\int_{M}\operatorname{tr}(u_{\infty}(t_{0}))\frac{\omega^{n}}{n!}=0$ and
(3.55) $\|u_{\infty}(t_{0})\|_{L^{\infty}(X\setminus\Sigma,H(t_{0}))}\leq r^{2}C_{3}.$
Furthermore, if $\Upsilon:R\times R\rightarrow R$ is a positive smooth function such that $\Upsilon(\lambda_{1},\lambda_{2})<(\lambda_{1}-\lambda_{2})^{-1}$ whenever $\lambda_{1}>\lambda_{2}$, then
(3.56) $\begin{split}&\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}(t_{0})\sqrt{-1}\Lambda_{\omega}(F_{H(t_{0}),\theta}))+\langle\Upsilon(u_{\infty}(t_{0}))(\overline{\partial}_{\theta}u_{\infty}(t_{0})),\overline{\partial}_{\theta}u_{\infty}(t_{0})\rangle_{H(t_{0})}\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.\end{split}$
By (3.55) and (3.14), we known that $\|\Lambda_{\omega}(F_{H(t_{0}),\theta})\|_{L^{1}(X\setminus\Sigma,H(t_{0}))}$ and $\|u_{\infty}(t_{0})\|_{L^{\infty}(X\setminus\Sigma,H(t_{0}))}$ are uniformly bounded (independent of $t_{0}$). Then, (3.56) implies there exists a uniform constant $\check{C}$ independent of $t_{0}$ such that $\int_{X\setminus\Sigma}|\overline{\partial}_{\theta}u_{\infty}(t_{0})|_{H(t_{0})}^{2}\frac{\omega^{n}}{n!}\leq\check{C}.$ By (3.32), we know that $\hat{H}$ and $H(t_{0})$ are locally mutually bounded each other. By choosing a subsequence, we have $)u_{\infty}(t_{0})\rightarrow u_{\infty}\in L_{1}^{2}(\Gamma(S_{\hat{H}}$)) weakly in local $L_{1}^{2}$ on $X\setminus\Sigma$ as $t_{0}\rightarrow 0$, where $u_{\infty}$ satisfies
(3.57) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty})\frac{\omega^{n}}{n!}=0,\quad and\quad\|u_{\infty}\|_{L^{1}(X\setminus\Sigma,\hat{H})}=1.$
Furthermore, we have
(3.58) $\lim_{t_{0}\rightarrow 0}\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}(t_{0})\sqrt{-1}\Lambda_{\omega}F_{H(t_{0}),\theta_{X_{\mathrm{reg}}}})\frac{\omega^{n}}{n!}=\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta_{X_{\mathrm{reg}}}})\frac{\omega^{n}}{n!}.$
Set $\hat{u}_{\infty}(t_{0})=(h(t_{0}))^{\frac{1}{2}}\cdot u_{\infty}(t_{0})\cdot(h(t_{0}))^{-\frac{1}{2}}.$ It is easy to check that: $\hat{u}_{\infty}(t_{0})\in S_{\hat{H}}(\mathcal{E}|_{X\setminus\Sigma})$ and $|\hat{u}_{\infty}(t_{0})|_{\hat{H}}=|u_{\infty}(t_{0})|_{H(t_{0})}$. Following the same argument in [73, Lemma 4.2], we have:
(3.59) $\lim_{t_{0}\rightarrow 0}\int_{\Omega}|\langle\tilde{\Upsilon}(u_{\infty}(t_{0}))(\overline{\partial}_{\theta}u_{\infty}(t_{0})),\overline{\partial}_{\theta}u_{\infty}(t_{0})\rangle_{H(t_{0})}-\langle\tilde{\Upsilon}(\hat{u}_{\infty}(t_{0}))(\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})),\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})\rangle_{\hat{H}}|\frac{\omega^{n}}{n!}=0,$
for any compact domain $\Omega\subset X\setminus\Sigma$ and any positive smooth function $\tilde{\Upsilon}:R\times R\rightarrow R$.
By (3.56), (3.58) and (3.59), for any given compact domain $\Omega\subset X\setminus\Sigma$ and any given positive number $\tilde{\epsilon}>0$, we have
(3.60) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})\frac{\omega^{n}}{n!}+\int_{\Omega}\langle\Upsilon(\hat{u}_{\infty}(t_{0}))(\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})),\overline{\partial}_{\theta}\hat{u}_{\infty}(t_{0})\rangle_{\hat{H}}\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}+\tilde{\epsilon}$
for small $t_{0}$. As we know that $\hat{u}_{\infty}(t_{0})\rightarrow u_{\infty}$ in $L^{2}(\Omega)$ as $t_{0}\rightarrow 0$, $|\hat{u}_{\infty}(t_{0})|_{\hat{H}}$ is uniformly bounded in $L^{\infty}$ and $|\overline{\partial}_{\phi}\hat{u}_{\infty}(t_{0})|_{\hat{H},\omega}$ is uniformly bounded in $L^{2}(\Omega)$. By the same argument as that in Simpson’s paper (Lemma 5.4 in [97]), we have
(3.61) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})\frac{\omega^{n}}{n!}+\|\Upsilon^{\frac{1}{2}}(u_{\infty})(\overline{\partial}_{\theta}u_{\infty})\|_{L^{q}(\Omega)}^{2}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}+2\tilde{\epsilon}$
for any $q<2$ and any $\tilde{\epsilon}>0$. Since $\tilde{\epsilon}$, $q<2$ and $\Omega$ are arbitrary, we get
(3.62) $\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})+\langle\Upsilon(u_{\infty})(\overline{\partial}_{\theta}u_{\infty}),\overline{\partial}_{\theta}u_{\infty}\rangle_{\hat{H}}\frac{\omega^{n}}{n!}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.$
By the above inequality and the Lemma 5.5 in [97], we can see that the eigenvalues of $u_{\infty}$ are constant almost everywhere. Let $\mu_{1}<\dots<\mu_{l}$ denote the distinct eigenvalue of $u_{\infty}$. Since $\int_{X\setminus\Sigma}\operatorname{tr}u_{\infty}\frac{\omega^{n}}{n!}=0$ and $\|u_{\infty}\|_{L^{1}}=1$, we must have $l\geq 2$. By Uhlenbeck and Yau’s regularity theorem of $L_{1}^{2}$-subbundle ([104]) and Simpson’s result (page 887 in [97]), $\pi_{\alpha}$ represents a saturated $\theta_{\mathrm{reg}}$-invarient coherent sub-sheaf $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ on the open set $X\setminus\Sigma$. Since $\mathcal{E}_{X}$ is reflexive, the trivial extension of $\mathcal{E}_{X_{\mathrm{reg}},\alpha}$ is a saturated subsheaf of $\mathcal{E}_{X}$ that $\theta_{X_{\mathrm{reg}}}$-invariant on $X_{\mathrm{reg}}$ by Lemma 2.5, which will be denoted by $\mathcal{E}_{\alpha}$. Let $\mathcal{E}_{\alpha,\mathrm{orb}}$ be the saturation of $\{\operatorname{image}(f_{\mu}^{*}\mathcal{E}_{\alpha}\xrightarrow{f_{\mu}^{*}}f_{\mu}^{*}\mathcal{E}_{X})\cap E_{\mu}\}$ on $E_{\mathrm{orb}}$. Recall Proposition 3.4, $f_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$ and thus there exists an analytic subspace $Z\subset X$ of codimension $\geq 2$ such that $\mathcal{E}_{\alpha,\mathrm{orb}}:=\{f_{\mu}^{*}\mathcal{E}_{\alpha}/\mathrm{torsion}\}$. It follows from Lemma 2.18, Lemma 2.17 and the Chern-Weil formula in the context of orbifolds (cf. [110, Section 2.2.2]) that
$\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{\alpha})=\deg_{f_{\mathrm{orb}}^{*}\omega}(\mathcal{E}_{\alpha,\mathrm{orb}})$
$=$ $\int_{Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma)}\operatorname{tr}(\pi_{\alpha_{\mathrm{orb}}}\sqrt{-1}\Lambda_{f_{\mathrm{orb}}^{*}\omega}F_{\hat{H}_{\mathrm{orb}},\theta_{\mathrm{orb}}})-|\overline{\partial}_{\theta_{\mathrm{orb}}}\pi_{\alpha,\mathrm{orb}}|_{\hat{H}_{\mathrm{orb}},f^{*}\omega_{\mathrm{orb}}}^{2}\frac{f^{*}\omega_{\mathrm{orb}}^{n}}{n!}$
$=$ $\int_{X\setminus\Sigma}\operatorname{tr}(\pi_{\alpha}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta_{X_{\mathrm{reg}}}})-|\overline{\partial}_{\theta_{X_{\mathrm{reg}}}}\pi_{\alpha}|_{\hat{H},\omega}^{2}\frac{\omega^{n}}{n!}$
where $\pi_{\alpha,\mathrm{orb}}$ denotes the weakly holomorphic subbundle associated to $\mathcal{E}_{\alpha,\mathrm{orb}}$. Furthermore, by the argument in Simpson’s paper (page 888 in [97]) and the inequality (3.62), we have
(3.63) $\begin{split}\int_{X\setminus\Sigma}\operatorname{tr}(u_{\infty}\sqrt{-1}\Lambda_{\omega}F_{\hat{H},\theta})+\langle\sum_{\alpha=1}^{l-1}(\mu_{\alpha+1}-\mu_{\alpha})(dP_{\alpha})^{2}(u_{\infty})(\overline{\partial}_{\theta}u_{\infty}),\overline{\partial}_{\theta}u_{\infty}\rangle_{\hat{H}}\leq-r^{-\frac{1}{2}}\frac{C^{\ast}}{\hat{C}_{1}}.\end{split}$
This complete the proof of (ii). ∎
3.3.3. Proof of Theorem 3.5
Let $H(t)$ be the long time solution of (3.9) on the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathrm{reg}})|_{X\setminus\Sigma}$ with initial metric $\hat{H}$. By Proposition 3.22 (ii), if the Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then $\|\log(H^{-1}(1)H(t))\|_{L^{1}}$ are uniformly bounded, i.e. we obtain uniform $C^{0}$-estimates on $X\setminus\Sigma$. By the standard parabolic estimates, we can get uniform local $C^{\infty}$-estimates, and then, by choosing a subsequence, $H(t)\rightarrow H_{\infty}$ in $C^{\infty}_{loc}(X\setminus\Sigma)$-topology, and (3.42) implies
(3.64) $\sqrt{-1}\Lambda_{\omega}(F_{H_{\infty}}+[\theta_{X_{\mathrm{reg}}},(\theta_{X_{\mathrm{reg}}})^{*H_{\infty}}])=\lambda\cdot\mathrm{Id}_{E}.$
Using Proposition 3.22 again, we known that, if the Higgs bundle $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus\Sigma}$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable, then
(3.65) $\lim_{t\rightarrow\infty}\int_{X\setminus\Sigma}|\sqrt{-1}\Lambda_{\omega}F_{H(t),\theta_{X_{\mathrm{reg}}}}-\lambda\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}|_{H(t)}^{2}\frac{\omega^{n}}{n!}=0.$
Combining (3.15), we obtain (3.11). This completes the proof of Theorem 3.5.
4. Characterization of harmonic bundles over the regular loci
This section begins with a review of basic facts (Section 4.1). It continues in Section 4.2 with the proof of part (1) of Theorem 1.13 and of the descent result for Higgs bundles from harmonic bundles (Proposition 4.11). The proof of Theorem 1.13 is concluded in Section 4.3.
4.1. Basic facts
We recall the following elementary facts about resolutions of singularities and maximally quasi-étale covers for later use.
Lemma 4.1 (cf. [38, Lemma 7.8] and [101]).
Let $\pi:\widehat{X}\rightarrow X$ be a resolution of singularities of a compact Kähler klt space $X$. The following hold:
  1. (4.1.1)
    For any locally free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ such that $\pi^{*}(\mathcal{E}_{X},\theta_{X})$ is semistable with (1.1) for some single Kähler form $\omega_{\widehat{X}}$, then $(\mathcal{E}_{X},\theta_{X})$ is semi-stable with respect to any nef forms $\eta_{0},\cdots,\eta_{n-2}$ and all Chern classes $c_{i}(\mathcal{E}_{X})\in H^{2i}(X,\mathbb{Q})$ vanish.
  2. (4.1.2)
    $\pi_{*}:\pi_{1}(\widehat{X})\rightarrow\pi_{1}(X)$ is an isomorphism. In particular, the pullback $\pi^{*}$ is a one-to-one map from $\mathrm{LSys}_{X}$ to $\mathrm{LSys}_{\widehat{X}}$ (resp. from $\mathrm{sLSys}_{X}$ to $\mathrm{sLSys}_{\widehat{X}}$).
Proof.
The item (4.1.1) follows from the same proof of [38, Lemma 7.8]. The item (4.1.2) follows from Takayama’s result [101] that the induced morphism of $\pi_{*}:\pi_{1}(\widehat{X})\rightarrow\pi_{1}(X)$ is an isomorphism, the proof is purely analytic and thus also holds for analytic klt spaces. ∎
Let us follow [34, 15, 55, 28] to review the concepts of maximally quasi-étale cover, which takes a central role in our construction.
Definition 4.2 (Maximally quasi-étale, [15, Definition 5.3]).
A normal complex space $X$ is said to be maximally quasi-étale if $X$ satisfies the following equivalent conditions:
  • (1)
    Any quasi-étale cover of $X$ is étale.
  • (3)
    Any étale cover of $X_{\mathrm{reg}}$ can be extended to a étale cover of $X$.
  • (3)
    The morphism of étale fundamental groups $\widehat{\pi}_{1}(X_{\mathrm{reg}})\rightarrow\widehat{\pi}_{1}(X)$ induced by the natural inclusion $i:X_{\mathrm{reg}}\hookrightarrow X$ is an isomorphism.
Remark 4.3 (The extension of linear representations).
If $X$ is maximally quasi-étale, then it follows from Malcev’s theorem that any linear representation of $\pi_{1}(X_{\mathrm{reg}})$ factors through $\pi_{1}(X)$ (cf. [36, Proof of Proposition 3.10]), i.e.
$\pi_{1}(X_{\mathrm{reg}})$$\pi_{1}(X)$$\mathrm{GL}(r,\mathbb{C}).$$\scriptstyle{i_{*}}$$\scriptstyle{\rho_{0}}$$\scriptstyle{\rho}$
Since $i_{*}:\pi_{1}(X_{\mathrm{reg}})\rightarrow\pi_{1}(X)$ is surjective (cf. [65, Proposition 2.10]), $\rho_{0}$ is semi-simple if and only if $\rho$ is semi-simple. This is equivalent to that, any (resp. semi-simple) local system on $X\setminus\Sigma$ for some analytic subspace of $\mathrm{codim}_{Z}X\geq 2$, can be extended to a (resp. semi-simple) local system on $X$ by combining $\pi_{1}(X_{\mathrm{reg}}\setminus\Sigma)\cong\pi_{1}(X_{\mathrm{reg}})$.
Proposition 4.4 (The existence of maximally quasi-étale covers).
Let $X$ be a compact complex klt space. Then there exists a compact klt space $Y$ and a quasi-étale Galois morphism $\gamma:Y\rightarrow X$ such that $Y$ is maximally quasi-étale.
Remark 4.5.
The existence of a maximally quasi-étale cover of a projective klt variety was proved by [34]. Building on the recent work of Fujino [29] of relative MMP in the context of analytic spaces, the existence of maximally quasi-étale covers for analytic klt spaces can be obtained by following the same method in [34] (see e.g. [28, Theorem 2.4 and Theorem 2.5]).
Lemma 4.6.
Let $X$ be a compact Kähler klt space and $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$ induced by some $G\subset\mathrm{Aut}(Y)$. Then the following hold:
  1. (4.6.1)
    For any Kähler class $\omega$ on $X$, $\gamma^{*}\omega$ is Kähler.
  2. (4.6.2)
    $\gamma$ is finite and étale over $X_{\mathrm{reg}}$. In particular, $\gamma^{-1}(X_{\mathrm{reg}})\subset Y_{\mathrm{reg}}$.
  3. (4.6.3)
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{{X_{\mathrm{reg}}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$ if and only if $(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})=\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{Y_{\mathrm{reg}}}$.
  4. (4.6.4)
    For any $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$ (resp. $\mathrm{sLSys}$), $(\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})})^{*}E_{X_{\mathrm{reg}}}$ extends to a $G$-equivariant local system $\rho_{Y}\in\mathrm{LSys}_{Y}$ (resp. $\mathrm{sLSys}_{Y}$). Conversely, the restriction of a $G$-equivariant local system $\rho_{Y}\in\mathrm{LSys}_{Y}$ (resp. $\mathrm{sLSys}_{Y}$) descends to a local system $\rho_{X_{\mathrm{reg}}}\in\mathrm{sLSys}_{X_{\mathrm{reg}}}$.
Proof.
The item (4.6.1) can be refered to [31, Proposition 3.6]. The item (4.6.2) follows from the purity of branch theorem due to Grauert-Remmert [32]. The item (4.6.3) follows from Lemma 2.21, Lemma 2.24 and the item (4.6.1). The item (4.6.4) follows from Remark 4.3. It also hold for semi-simple local systems bacause $\gamma_{*}:\pi_{1}(\gamma^{-1}(X_{\mathrm{reg}}))\subset\pi_{1}(Y_{\mathrm{reg}})$ is injective by the item (4.6.2). ∎
4.2. Harmonic bundles and descent result
This section is a study on harmonic bundle structure over the regular locus of a compact Kähler klt space $X$. The following standard definition is given to fix the conventions (cf. [106, Section 2]).
Let $(E,D)$ be a flat bundle equipped with a Hermitian metric $H$ on a complex complex manifold $M$. There exists a unique decomposition of $D$ by
(4.1) $D=D_{H}+\Psi_{H}$
where $D_{H}$ is a $H$-connection and $\Psi_{H}\in\Lambda^{1}(M,\operatorname{End}(E))$ is self-adjoint with respect to $H$. We define
(4.2) $G_{H}:=(D_{H}^{0,1}+\Psi_{H}^{1,0})^{2}=(D_{H}^{0,1})^{2}+D_{H}^{0,1}\Psi_{H}^{1,0}+\Psi_{H}^{1,0}\wedge\psi_{H}^{1,0}.$
Definition 4.7 (Harmonic bundle).
We say that $(E,D,H)$ is a harmonic bundle or a flat bundle $(E,D)$ admits a pluri-harmonic metric $H$ if $G_{H}=0$, which is equivalent to that
(4.3) $(D_{H}^{0,1})^{2}=0,\ D_{H}^{0,1}\Psi_{H}^{1,0}=0\text{ and }\Psi_{H}^{1,0}\wedge\psi_{H}^{1,0}=0.$
Thus, $(E,D,H)$ is a harmonic bundle if and only if $(E,\bar{\partial},\theta):=(E,D_{H}^{0,1},\Psi_{H}^{1,0})$ is a Higgs bundle and we could write the decomposition (4.1) as
(4.4) $D=\partial+\bar{\partial}+\theta+\theta^{*},$
4.2.1. Pluriharmonic metrics on Harmonic bundles
This subsubsection is devoted to prove the following statement.
Proposition 4.8.
Let $X$ be a compact Kähler klt space and $Z$ be an analytic subspace of $X$ with $\mathrm{codim}_{Z}X\geq 2$. Suppose that $X\setminus Z$ is smooth. Then the following hold:
  • (1)
    A flat bundle on $X\setminus Z$ is semi-simple if and only if it admits a pluri-harmonic metric, which is unique up to some constant on each simple component.
  • (2)
    Any harmonic bundle $(E_{X\setminus Z},D_{X\setminus Z},H_{X\setminus Z})$ on $X\setminus Z$ could extend to a harmonic bundle $(E_{X_{\mathrm{reg}}},D_{X_{\mathrm{reg}}},H_{X_{\mathrm{reg}}})$ on $X_{\mathrm{reg}}$.
Because $i_{*}:\pi_{1}(X_{\mathrm{reg}}\setminus Z)\rightarrow\pi_{1}(X_{\mathrm{reg}})$ is an isomorphism, any semi-simple flat bundle on $X\setminus Z$ could extends to a semi-simple flat bundle on $X_{\mathrm{reg}}$ and then (2) is a direct consequence of (1).
Remark 4.9.
When $Z=\emptyset$, Proposition 4.8 rests on the work of Corlette and Donaldson on harmonic metrics [17, 25]; see [16, 18, 60, 61, 70, 76, 98, 83, 106] for various generalizations.
Proof of Proposition 4.8.
[…]
It remains to verify the uniqueness and that the existence of a pluri-harmonic metric implies semi-simpleness. The classical proof in the non-compact context via the integral method is not available because we do not put any assumption such as “admissibility” on pluri-harmonic metrics. We adapt another method via the maximum principle, which is well‑adapted to our setting thanks to Grauert–Remmert’s extension theorem [32].
Lemma 4.10.
Let $(E,D,H)$ be a harmonic bundle on a complex manifold $M=X\setminus Z$ for some compact analytic space $X$ and an analytic subspace $Z$ of codimension at least $2$. Then for any $D$-invariant subbundle $W$ of $E$ such that $(W,D_{W})$ is simple where $D_{W}=D|_{W}$, it holds that
(4.5) $(E,D,H)=(W,D_{W},H_{W})\oplus(W^{\perp},D_{W^{\perp}},H_{W^{\perp}})$
splits as harmonic bundles, where $H_{W}=H|_{W}$, $W^{\perp}$ is the orthogonal complement of $W$ with respect to $H$ and $D_{W^{\perp}}$ (resp. $H_{W^{\perp}}$) is the induced connection (resp. the induced Hermitian metric) on $W^{\perp}$.
Proof of Lemma 4.10.
$D$ takes the form
(4.6) $D=\left(\begin{matrix}D_{W}&\beta_{W}\\ 0&D_{W^{\perp}}\end{matrix}\right)$
where $\beta_{W}\in\Lambda^{1}(M,\operatorname{Hom}(W^{\perp},W))$. Recall (4.4) that $D$ uniquely decomposes as $D_{H}+\Psi_{H}$. It is easily known that $D_{H}$ and $\Psi_{H}$ take the form
(4.7) $D_{H}=\left(\begin{matrix}D_{H_{W}}&\frac{1}{2}\beta_{W}\\ -\frac{1}{2}\beta^{*}_{W}&D_{H_{W^{\perp}}}\end{matrix}\right)\ \text{ and }\ \Psi_{H}=\left(\begin{matrix}\Psi_{H_{W}}&\frac{1}{2}\beta_{W}\\ \frac{1}{2}\beta^{*}_{W}&\Psi_{H_{W^{\perp}}}\end{matrix}\right)$
where we decompose $D_{W}=D_{H_{W}}+\Psi_{H_{W}}$ and $D_{W^{\perp}}=D_{H_{W^{\perp}}}+\Psi_{H_{W^{\perp}}}$ (cf. [106, Section 2.3]). Recall the pluri-harmonicity of $H$ means (4.3). Combining this with (4.7) gives
(4.8) $D_{H_{W}}^{0,1}\Psi_{H_{W}}^{1,0}-\frac{1}{4}\beta_{W}^{1,0}\wedge(\beta_{W}^{1,0})^{*}+\frac{1}{4}\beta_{W}^{0,1}\wedge(\beta_{W}^{0,1})^{*}=0.$
Recall the simple flat bundle $(W,D_{W})$ admits a pluri-harmonic metric $K$ as explained above. Then $D_{K}^{0,1}\Psi_{K}^{1,0}=0$ and (4.8) imply
(4.9) $\begin{split}\sqrt{-1}\partial\bar{\partial}\log\det(h)=&2\sqrt{-1}\operatorname{tr}(D_{H_{W}}^{0,1}\Psi_{H_{W}}^{1,0}-D_{K}^{0,1}\Psi_{K}^{1,0})\\ =&\frac{1}{2}\sqrt{-1}\operatorname{tr}\left(\beta_{W}^{1,0}\wedge(\beta_{W}^{1,0})^{*}-\beta_{W}^{0,1}\wedge(\beta_{W}^{0,1})^{*}\right)\geq 0,\end{split}$
where $h=K^{-1}H_{W}$ (cf. [93, (2.25)] for the derivation of the first equality). Consequently, $\log\det(h)$ is a plurisubharmonic function on $M$ and so can be extended to a plurisubharmonic function on $X$ due to the extension theorem by Grauert-Remmert [32, page 181] since $\mathrm{codim}_{Z}X\geq 2$. Thus, $\log\det(h)$ must be constant. Again using (4.9), we deduce that $\beta_{W}^{1,0}=\beta_{W}^{0,1}=0$ and so $\beta_{W}=0$, which completes the proof. ∎
To prove the uniqueness, it suffices to consider the case when $(E,D)$ is simple by Lemma 4.10. Suppose that $H,K$ are two pluri-harmonic metrics on $(E,D)$, then we have
(4.10) $\sqrt{-1}\partial\bar{\partial}\operatorname{tr}(h)=\frac{1}{2}\sqrt{-1}\{Dh\cdot h^{-\frac{1}{2}},\overline{Dh\cdot h^{-\frac{1}{2}}}\}_{K}+\operatorname{tr}(h\sqrt{-1}(G_{H}-G_{K})),$
where $h=K^{-1}H\in C^{\infty}(M,\operatorname{End}(E))$ (cf. [93, (2.36)]). Consequently, $\operatorname{tr}(h)$ is a global plurisubharmonic function on $M$ because $G_{H}=0$ and $G_{K}=0$ (cf. (4.3)) and so must be some constant as above. Applying (4.10) again gives $Dh=0$ over $M$. Then $h=c\operatorname{Id}_{E}$ for some constant $c>0$. Otherwise, $(E,D)$ must split according to the eigenspaces of $h$. Therefore $H=cK$.
It remains to show that the existence of a pluri-harmonic metric $H$ implies the semi-simpleness of $(V,D)$, which can be concluded by Lemma 4.10 and an induction argument because any flat bundle $(E,D)$ includes a $D$-invariant simple subbundle. Thus, the proof of Proposition 4.8 is complete. ∎
4.2.2. Descent results
One of the key technical components of showing Theorem 1.13 (2) is the following descent result.
Proposition 4.11.
[…]
Thanks to Daniel’s work [19] on introducing the category of variations of loop Hodge structures [19], we are able to study harmonic bundles using the classical tools of Hodge theory, notably the existence of a period map, which record the information of the structure of the induced Higgs sheaf as in the following statement.
Lemma 4.12.
Let $(E,D,h)$ be a harmonic bundle over a simply connected complex manifold $X$ and $f:M\rightarrow\mathcal{D}$ be the induced period map introduced in [19, Section 3.2]. Write $D=\partial+\bar{\partial}+\theta+\theta^{*}$ by the decomposition (4.4) and set $\mathcal{E}_{X}:=(E,\bar{\partial})$, then there exists a locally free coherent sheaf $\mathcal{E}_{D}$ on $\mathcal{D}$ such that $\mathcal{E}_{X}:=f^{*}\mathcal{E}_{D}$.
To see this, it’s necessary to outline Daniel’s construction of the period map. The reader is encouraged to have a look on [19] for a concise definition when reading the following construction because involving concepts not present in the classic case.
Construction 4.13 (The period map, cf. [19, Section 3]).
Let $x_{0}\in x$ be a base point and $\mathcal{H}=L^{2}(S^{1},\operatorname{End}(E))$, then the fiber of the canonical outgoing Krein bundle $(\mathcal{H},\mathcal{B},\mathcal{T})$ induced by $(E,h)$ over $x_{0}$ can be identified with the canonical Krein space $(H:=L^{2}(S^{1},\mathbb{C}^{r}),B,T)$ by the identification of $(E,h)_{x_{0}}$ with the standard Hermitian inner product on $\mathbb{C}^{r}$ (see [19, Section 2.1.3]), where $r=\operatorname{rank}(E)$ and $S^{1}=\{z\in\mathbb{C}:|z|=1\}$.
When there is no risk of confusion, the trivial bundle $Y\times H$ over a complex manifold $Y$ will be simply denoted by $H_{Y}$. Set
$D_{\lambda}=\partial+\bar{\partial}+\lambda^{-1}\theta+\lambda\theta^{*}$
for any $\lambda\in S^{1}$. Then $\{D_{\lambda}\}$ can be viewed a flat connection $\widetilde{D}$ on $\mathcal{H}$; moreover, $\mathcal{B}$ and $\mathcal{T}$ are $\widetilde{D}-$flat (see [19, Appendix A]). In particular, the canonical Krein space with flat connection $\widetilde{D}$ has the trivialization
$(\mathcal{H},\mathcal{B},\mathcal{T},\widetilde{D})\cong(H_{X},B_{X},T_{X},d_{X}),$
where $B_{X}$ (resp. $T_{X}$) is the trivial Krein metric (resp. the trivial outgoing operator) by identifying with the base point and $d_{X}$ is the tribial connection.
Using Fourier series decomposition, any element $f\in\mathcal{H}$ can be written as $s(\lambda)=\sum\limits_{n\in\mathbb{Z}}s_{n}\lambda^{n}$ such that $(s_{n}(x))_{n\in\mathbb{Z}}\in l^{2}(C^{r})$ for any $x\in X$. The canonical outgoing subbundle is given by $\mathcal{W}=L^{2}_{+}(S^{1},\operatorname{End}(E))$ (cf. [19, Section 2.1.3]), the set of all Fourier series with non-negative degrees. In particular, for any $x$, $\mathcal{W}_{x}$ can be identified with an outgoing subspace $W$ of $H$, which varies when $x$ varies over $X$. In summary, we have the following isomorphism of variations of loop Hodge structures
(4.11) $\gamma:(\mathcal{H},\mathcal{B},\mathcal{T},\mathcal{W},\widetilde{D})\cong(H_{X},B_{X},T_{X},W_{X},d_{X}),$
where $(H_{X},B_{X},T_{X})$ is trivial with respect to $d_{X}$ while $W$ is not.
Denote $\mathrm{Gr}(H)$ be the outgoing Grassmannian of $H$, that is the set of all outgoing subspaces of $H$. By [19, Proposition 3.7 and Remark 3.8], we can identify $\mathrm{Gr}(H)$ with $\Lambda_{\sigma}^{\infty}G/K$ (cf. [19, p. 619] for its definition), where $G=\mathrm{GL}(r,\mathbb{C})$ and $K=\mathrm{U}(r,C)$. Therefore, $f:x\mapsto W_{x}$ gives a map from $X$ to $\Lambda_{\sigma}^{\infty}G/K$. Denote $\mathcal{D}=\Lambda_{\sigma}G/K$ be the period domain defined in [19, Section 3.1.4]. $\mathcal{D}$ is an infinite-dimensional complex Hilbert manifold and $f(x)$ indeed lies in $\mathcal{D}$ (cf. [19, Lemma 3.19]). Then $f:X\rightarrow\mathcal{D}$ is defined as the period map of $(E,D,h)$.
Proof of Lemma 4.12.
[…]
(4.12) $(E,\bar{\partial})\cong(\mathcal{W}/\mathcal{T}\mathcal{W},\widetilde{D}^{0,1})\cong(W_{X}/T_{X}W_{X},d^{0,1})$
where the second isomorphism is given by (4.11).
$\mathrm{Gr}(H)=\Lambda_{\sigma}^{\infty}G/K$ implies that there exists a tautological bundle over $\mathcal{D}$ defined by $(H^{0,\mathcal{D}})_{W\in\mathcal{D}}$, where $H^{0,\mathcal{D}}$ is the orthogonal complement of $TW$ in $W$ (cf. [19, Definition 3.12]). Let $d_{\mathcal{D}}$ be the trivial connection of the trivial bundle $H_{\mathcal{D}}$, then it induces a holomorphic structure $\bar{\partial}_{\mathcal{D}}$ on $H^{0,\mathcal{D}}$. According to the construction of the period map $f$ explained as above,
$(W_{X}/T_{X}W_{X},d_{X}^{0,1})=f^{*}(H^{0,D},d_{\mathcal{D}}^{0,1})$
since $f$ is holomorphic by [19, Theorem 3.18]. Via the identification (4.12), we conclude that $\mathcal{E}_{X}=(E,\bar{\partial})=f^{*}(H^{0,D},d_{\mathcal{D}}^{0,1})$. Taking $\mathcal{E}_{\mathcal{D}}=(H^{0,D},d_{\mathcal{D}}^{0,1})$ and the proof is complete. ∎
Building on Lemma 4.12 and [19, Theorem 1.2], we can adapt the arguments of [37, Corollary 5.8] to prove Proposition 4.11, where treats Higgs bundles from variations of Hodge structures.
Proof of Proposition 4.11.
Recall Takayama’s result [101] that for any $x\in X$, there exists an open neighborhood $U$ of $x$ such that $M:=\pi^{-1}(U)$ is simply connected. We shall show that there exists a locally free sheaf $\mathcal{E}_{U}$ on $U$ such that $\mathcal{E}_{\widehat{X}}|_{\pi^{-1}(U)}=\pi^{*}\mathcal{E}_{U}$, which suffices to conclude the desired statement by Lemma 2.4.
Let $f:\pi^{-1}(U)\rightarrow\mathcal{D}$ be the induced period map of the harmonic bundle $(E,D,h)|_{\pi^{-1}(U)}$. Building on Lemma 4.12, there exists a locally free sheaf $\mathcal{E}_{\mathcal{D}}$ such that $\mathcal{E}_{X}|_{\pi^{-1}(U)}=f^{*}\mathcal{E}_{\mathcal{D}}$. If $f$ factories through $\pi$, i.e., $f=g\circ\pi$ for some holomorphic map $g:U\rightarrow\mathcal{D}$, then $\mathcal{E}_{X}=\pi^{*}(g^{*}\mathcal{E}_{\mathcal{D}})$. It suffices to show that $f$ is constant on $\pi^{-1}(x)$ for any $x\in U$. Fix a point $x\in U$. For any rational curve $\gamma:\mathbb{P}^{1}\rightarrow\pi^{-1}(x)$, $f\circ\gamma$ is again holomorphic and horizontal (cf the definition in [19, 3.1.6]) and so is a period map induced by a harmonic bundle structure over $\mathbb{P}^{1}$ (cf. [19, Theorem 1.2]). However, any harmonic bundle structure over $\mathbb{P}^{1}$ is a trivial bundle and so $x\mapsto W_{x}$ in (4.11) is constant, which implies that $f\circ\gamma$ is a constant map. Recall that $\pi^{-1}(x)$ is rationally chain connected (cf. [30]), i.e., any two points of $\pi^{-1}(x)$ can be connected by a chain of rational curves. We conclude that $f$ is constant on $\pi^{-1}(x)$ for any $x\in U$ and the proof is complete. ∎
4.3. Proof of Theorem 1.13
Let us complete the proof of Theorem 1.13. Let $X$ be a compact Kähler klt space and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{pHiggs}_{X_{\mathrm{reg}}}$. Building on Theorem 3.2, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is locally free and admits a pluri-harmonic metric $H_{X\setminus Z}$ outside an analytic subspace $Z$ such that $\mathrm{codim}_{X}Z\geq 2$ and $X\setminus Z$ is smooth. It follows from of Proposition 4.8 that the associated harmonic bundle structure of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus Z}$ extends to a harmonic bundle on $X_{\mathrm{reg}}$ and so $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})|_{X\setminus Z}$ extends to a locally free Higgs sheaf, which coincides with $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ by the reflexivity.
Thus we have verified the “only if part” of Theorem 1.13 (2); moreover, Theorem 1.13 (1) directly follows from Proposition 4.8 by substituting $X\setminus Z=X_{\mathrm{reg}}$. The remaining statements of Theorem 1.13 are included by the following statement.
Proposition 4.14.
Let $X$ be a compact Kähler klt space space. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a locally free Higgs sheaf equipped with a pluri-harmonic metric $H_{\mathrm{reg}}$ on $X_{\mathrm{reg}}$. Then for any maximally quasi-étale cover $\gamma:Y\rightarrow X$ induced by some $G\subset\mathrm{Aut}(Y)$, the following hold:
  1. ($4.1$)
    $\gamma^{[*]}\mathcal{E}_{X}$ is locally free and in particular $\gamma^{[*]}\theta_{{X_{\mathrm{reg}}}}$ can be extended to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}=\gamma^{[*]}\mathcal{E}_{X}$.
  2. (4.14.2)
    For any resolution of singularities $\pi:\widehat{Y}\rightarrow Y$, $\pi^{*}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{pHiggs}_{\widehat{Y}}$.
  3. (4.14.3)
    $(\mathcal{E}_{Y},\theta_{Y})$ is polystable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$.
In particular, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is polystable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ and the orbifold Chern classes of $\mathcal{E}_{X}$ vanish.
Proof.
[…]
(4.13) $\begin{split}&\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{F}_{Y})=\deg_{(\pi^{*}\omega_{0},\cdots,\pi^{*}\omega_{n-2})}(\mathcal{F}_{\widehat{Y}})\\ &=-\int_{\widehat{Y}\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}}\sqrt{-1}\operatorname{tr}(\partial\pi_{\mathcal{F}_{\widehat{Y}}}^{H_{\widehat{Y}}}\wedge\bar{\partial}\pi_{\mathcal{F}_{\widehat{Y}}}^{H_{\widehat{Y}}})\wedge\pi^{*}\omega_{0}\wedge\cdots\pi^{*}\omega_{n-2}\\ &\leq 0=\deg_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{Y})\end{split}$
since $F_{H_{\widehat{Y}},\theta_{\widehat{Y}}}=0$ and $c_{1}(\mathcal{E}_{Y})=0$ by (4.1.1), where $\Sigma_{\mathcal{F}_{\widehat{Y}}}$ is the singular set of $\mathcal{F}_{\widehat{Y}}$. Suppose that $(\mathcal{E}_{Y},\theta_{Y})$ is not $(\omega_{0},\cdots,\omega_{n-2})$-stable, then there exists some $\mathcal{F}_{Y}$ such that “=” in the last line of (4.13) holds, then $\partial\pi_{\widehat{Y}}^{H_{\widehat{Y}}}=0$ on $\pi^{-1}(Y_{\mathrm{reg}})\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}$ since $\pi^{*}\omega_{0},\cdots,\pi^{*}\omega_{n-2}$ are strictly positive on $\pi^{-1}(Y_{\mathrm{reg}})$. Thus
$(\mathcal{E}_{\widehat{Y}},\theta_{\widehat{Y}})=(\mathcal{F}_{\widehat{Y}},\theta_{\widehat{\mathcal{F}}})\oplus(\mathcal{F}_{\widehat{Y}}^{\perp},\theta_{\widehat{\mathcal{F}}^{\perp}})$
splits as locally free Higgs sheaves on $\pi^{-1}(Y_{\mathrm{reg}})\setminus\Sigma_{\mathcal{F}_{\widehat{Y}}}$, where $\theta_{\widehat{\mathcal{F}}}$ (resp. $\theta_{\widehat{\mathcal{F}}^{\perp}}$) is the induced Higgs field on $\mathcal{F}_{\widehat{Y}}$ (resp. $\mathcal{F}_{\widehat{Y}}^{\perp}$). Consequently,
$(\mathcal{E}_{{Y}},\theta_{{Y}})=(\mathcal{F}_{{Y}},\theta_{{\mathcal{F}}})\oplus(\mathcal{F}_{{Y}}^{\perp},\theta_{{\mathcal{F}}^{\perp}})$
splits as locally free Higgs sheaves on $Y_{\mathrm{reg}}\setminus\Sigma_{\mathcal{F}_{Y}}$, where $\mathcal{F}_{Y}^{\perp}$ is the orthogonal complement with respect to $\pi_{*}H_{\widehat{Y}}$ and $\Sigma_{\mathcal{F}_{Y}}$ is the singular set of $\mathcal{F}_{Y}$. Then the induced Hermitian metric $H_{\mathcal{F}_{Y}}$ (resp. $H_{\mathcal{F}_{Y}^{\perp}}$) on $\mathcal{F}_{Y}$ (resp. $\mathcal{F}_{Y}^{\perp}$) is pluri-harmonic. Since $\Sigma_{\mathcal{F}_{Y}}$ has codimension at least $2$, then $(\mathcal{F}_{Y},\theta_{\mathcal{F}_{Y}})$ and $(\mathcal{F}_{Y}^{\perp},\theta_{\mathcal{F}_{Y}^{\perp}})$ extend to a locally free Higgs sheaf that arises from a harmonic bundle structure by the same argument as above. Applying an induction argument, we verify the $(\omega_{0},\cdots,\omega_{n-2})$-polystability of $(\mathcal{E}_{Y},\theta_{Y})$.
Let us complete the proof. The item (4.14.2), the item (4.1.1), Remark 2.22 and Lemma 2.21 together implies that the orbifold Chern classes of $\mathcal{E}_{X}$ of $\mathcal{E}_{X_{\mathrm{reg}}}$ vanishes. On the other hand, the item ((4.14.3)) implies $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is stable with respect to $\gamma^{*}\omega_{0},\cdots,\gamma^{*}\omega_{n-2}$ for any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$ and in particular, is $G$-stable, which implies the $(\omega_{0},\cdots,\omega_{n-2})$-polystability of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ by Lemma 2.24. ∎
5. Non-abelian Hodge correspondence over Kähler klt spaces
This section first proves the two key ingredients outlined in Section 1.3: Theorem 1.12 (in Section 5.1) and Theorem 1.15 (in Section 5.2). We then prove Theorems 1.3 and 1.7 in Section 5.3, thereby establishing the nonabelian Hodge correspondences for compact Kähler klt spaces. Finally, we prove Corollary 1.9, Corollary 1.10 and Theorem 1.11 in Section 5.4.
5.1. Proof of Theorem 1.12
The proof of Theorem 1.12 will be divided into two parts. We review the following statements, which are useful in the proof of Theorem 1.12 and 1.15.
Lemma 5.1.
Let $0\rightarrow\mathcal{S}\rightarrow\mathcal{E}\rightarrow\mathcal{Q}\rightarrow 0$ be a short exact sequence of coherent sheaves on a Cohen-Macaulay, normal and irreducible complex space $X$ of dimension $n$. Suppose that $\mathcal{E}$ is reflexive, $\mathcal{S}$ and $\mathcal{Q}^{\vee\vee}$ is locally free. If addition $\mathcal{Q}$ is locally free in codimension $2$, then $\mathcal{E}$ and $\mathcal{Q}$ are locally free.
Lemma 5.1 follows from the following result that come back to Scheja and Trautmann [95, 103].
Proposition 5.2 (cf. [100, Pages 142]).
Let $X$ be a complex space, $A\subset X$ a closed analytic subset and $\mathcal{F}$ be a coherent sheaf on $X$. Fix a non-negative integer $k$. Then the following statements are equivalent:
  • (1)
    $\dim(A\cap S_{\mu+k+1}(\mathcal{F}))\leq\mu$ for all $\mu\in\mathbb{Z}$.
  • (2)
    for every open subset $U\subset X$, the restrictions $H^{i}(U,\mathcal{F})\rightarrow H^{i}(U\setminus A,\mathcal{F})$ are bijective for $i<k$ and injective for $i=k$.
We give an explanation on (Proposition 5.2 $\Rightarrow$ Lemma 5.1) for reader’s convenience. By the assumption of Lemma 5.1, there exists an analytic subspace $Z\subset X$ of codimension at least $3$ such that $\mathcal{Q}$ and $\mathcal{E}$ is locally free outside $Z$. Since $\mathcal{Q}^{\vee}=\mathcal{Q}^{\vee\vee\vee}$ and $\mathcal{S}$ are locally free, then for any $x\in X$,
[…]
where $\mathrm{codh}$ denotes the depth and third equality follows from the dimension formula. Thus the singularity set $S_{n-1}(\mathcal{S}\otimes\mathcal{Q}^{\vee}):=\{x\in X:\mathrm{codh}(\mathcal{S}\otimes\mathcal{Q}^{\vee})_{x}\leq n-1\}$ of $\mathcal{S}\otimes\mathcal{Q}^{\vee}$ is empty and $S_{n}(\mathcal{S}\otimes\mathcal{Q}^{\vee})=X$, we have that $\dim(A\cap S_{\mu+3})\leq\mu\ \text{ for any integer }\mu.$ Therefore, the restriction $H^{1}(X,\mathcal{S}\otimes\mathcal{Q}^{\vee})\rightarrow H^{1}(X\setminus Z,\mathcal{S}\otimes\mathcal{Q}^{\vee})$ is bijective by Proposition 5.2. This implies $\mathcal{F}|_{X\setminus Z}$ extends to a locally free Higgs sheaf $\mathcal{F}^{\prime}$ on $X$ that is the extension of $\mathcal{S}$ by $\mathcal{Q}^{\vee\vee}$. By the reflexivity, $\mathcal{F}$ coincides with $\mathcal{F}^{\prime}$ and the proof is complete.
Lemma 5.3 (The Hodge index Theorem, cf. [53, Lemma 3.17]).
Suppose that $\mathcal{Q}$ is a reflexive sheaf on $X$. Then for nef and big classes $\alpha_{0},\cdots,\alpha_{n-2}$, we have
(5.1) $\widehat{c}_{1}^{2}(\mathcal{Q})\cdot[\alpha_{1}]\cdots[\alpha_{n-2}]\leq\frac{(\widehat{c}_{1}(\mathcal{Q})\cdot[\alpha_{0}]\cdots[\alpha_{n-2}])^{2}}{[\alpha_{0}]^{2}\cdot[\alpha_{1}]\cdots[\alpha_{n-2}]}.$
The equality holds if and only if $\widehat{c}_{1}(\mathcal{E})\cdot[\gamma]\cdot[\alpha_{0}]\cdots[\alpha_{n-2}]=0$ for any $\gamma\in H^{2}(X,\mathbb{R})$.
5.1.1. Characterization of semistable reflexive Higgs sheaf by extension
We first prove part (1) of Theorem 1.12.
Theorem 5.4.
Let $X$ be a compact Kähler klt space of dimension $n$, $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$ and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ be a reflexive Higgs sheaf over $X_{\mathrm{reg}}$. Then the following statements are equivalent:
  1. (i)
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is semistable and satisfies (1.3) with respect to some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$.
  2. (ii)
    $\mathcal{E}_{Y}=\gamma^{[*]}\mathcal{E}_{X}$ is locally free, in particular $\gamma^{*}\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$ can be extended to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}$; $\mathcal{E}_{Y}$ admits a filtration of $\theta_{Y}$-invariant locally free sheaves
    $0=\mathcal{E}_{0}\subsetneq\mathcal{E}_{1}\subsetneq\cdots\subsetneq\mathcal{E}_{l}=\mathcal{E}_{Y}$
    such that each quotient $\mathcal{Q}_{k}:=\mathcal{E}_{k}/\mathcal{E}_{k-1}$ is locally free and its all Chern classes vanish; $\mathcal{Q}_{k}$ with the induced Higgs field $\theta_{k}$ is $(\gamma^{*}\omega_{0},\cdots,\gamma^{*}\omega_{n-2})$-stable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$.
  3. (iii)
    Orbifold Chern classes $\widehat{c}_{1}(\mathcal{E}_{X}),\widehat{c}_{1}^{2}(\mathcal{E}_{X}),\widehat{c}_{2}(\mathcal{E}_{X})$ of $\mathcal{E}_{X}$ vanish and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is semistable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$.
Ditto for polystability (in this case, the length $l$ of the filtration equals $1$).
Proof of Theorem 5.4.
The implication $(iii)\Rightarrow(i)$ is obvious and $(ii)\Rightarrow(iii)$ follows from Lemma 2.9, Lemma 2.24 and Lemma 2.21. It suffices to verify $(i)\Rightarrow(ii)$. According to the same reason of provining $(ii)\Rightarrow(iii)$, we may assume that $X$ is maximally quasi-étale and $\gamma$ is the identity map.
If $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then the proof is completed by applying Theorem 1.13, Proposition 4.14 and Lemma 4.1. Because any reflexive Higgs sheaf of $\operatorname{rank}1$ is $(\omega_{0},\dots,\omega_{n-2})$-stable, we could argue by induction on $\operatorname{rank}(\mathcal{E}_{X})$. Assuming that Theorem 5.4 holds for $\operatorname{rank}(\mathcal{E}_{X})<r$. Now let $\operatorname{rank}(\mathcal{E}_{X})=r$. Without loss of generality, we assume that $(\mathcal{E}_{X{\mathrm{reg}}},\theta_{\mathcal{E}_{X{\mathrm{reg}}}})$ is not $(\omega_{0},\dots,\omega_{n-2})$-stable.
The construction of $\mathcal{F}_{X}$ in this paragraph is standard, we include an explanation for reader’s convenience. It follows from Remark 2.6 that
$\mathcal{B}:=\{0\neq\mathcal{H}_{X}\subsetneq\mathcal{E}_{X}:\mathcal{H}_{X_{\mathrm{reg}}}\text{ is $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$-invariant and }\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{H}_{X})=\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})\}$
is nonempty. Taking $\mathcal{F}_{X}\in\mathcal{B}$ with the maximal rank. We may assume that $\mathcal{F}_{X}$ is saturated by Remark 2.6. In particular, the induced Higgs sheaf $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable. Set $\mathcal{Q}_{X_{\mathrm{reg}}}:=\mathcal{E}_{X_{\mathrm{reg}}}/\mathcal{F}_{X_{\mathrm{reg}}}$. Then a direct computation gives
$\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{Q}_{X})=\frac{(c_{1}(\mathcal{E}_{X})-c_{1}(\mathcal{F}_{X}))\cdot[\omega_{0}]\cdots[\omega_{n-2}]}{\operatorname{rank}(\mathcal{E}_{X})-\operatorname{rank}(\mathcal{F}_{X})}=\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})=0.$
The induced torsion-free Higgs sheaf $(\mathcal{Q}_{X_{\mathrm{reg}}},\theta_{\mathcal{S}_{X_{\mathrm{reg}}}})$ must be $(\omega_{0},\cdots,\omega_{n-2})$-stable. Otherwise, there exists a saturated subsheaf $\mathcal{G}_{X}\subsetneq\mathcal{Q}_{X}$ that is $\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}}$-invariant such that $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{G}_{X})\geq\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{Q}_{X}).$ Let $\mathcal{S}_{X}=\mathcal{G}_{X}+\mathcal{F}_{X}$, then we have the short exact sequence
$0\rightarrow\mathcal{G}_{X}\rightarrow\mathcal{S}_{X}\rightarrow\mathcal{F}_{X}\rightarrow 0,$
$\mathcal{S}_{X_{\mathrm{reg}}}$ is $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$-invariant and we have $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{S}_{X})\geq\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})$ by a direct computation, which contradicts that $\mathcal{F}_{X}$ has the maximal rank in $\mathcal{B}$.
It follows from Proposition 2.15, Bogomolov-Gieseker inequality (3.1) and Lemma 5.3 that
(5.2) $\begin{split}0=&\widehat{\mathrm{ch}}_{2}(\mathcal{E}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\widehat{\mathrm{ch}}_{2}(\mathcal{F}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]+\widehat{\mathrm{ch}}_{2}(\mathcal{Q}_{X}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\frac{1}{\operatorname{rank}(\mathcal{F}_{X})}\widehat{c}_{1}^{2}(\mathcal{F}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]+\frac{1}{\operatorname{rank}(\mathcal{Q}_{X})}\widehat{c}_{1}^{2}(\mathcal{Q}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\frac{1}{\operatorname{rank}\mathcal{F}_{X}}\frac{\big(\widehat{c}_{1}(\mathcal{F}_{X})\cdot[\omega_{0}]\cdots[\omega_{n-2}]\big)^{2}}{[\omega_{0}]\cdot[\omega_{0}]\cdots[\omega_{n-2}]}+\frac{1}{\operatorname{rank}\mathcal{F}_{X}}\frac{\big(\widehat{c}_{1}(\mathcal{Q}_{X}^{\vee\vee})\cdot[\omega_{0}]\cdots[\omega_{n-2}]\big)^{2}}{[\omega_{0}]\cdot[\omega_{0}]\cdots[\omega_{n-2}]}=0\end{split}$
since $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})=0$. Hence, each inequality is in fact an equality, which implies the following statements:
  • (i)
    $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ is a $(\omega_{0},\cdots,\omega_{n-2})$-semistable reflexive Higgs sheaf with (1.3); By the induction hypothesis, $\mathcal{F}_{X}$ is locally free and $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ satisfies property (2) of Theorem 5.4.
  • (ii)
    $(\mathcal{Q}_{X_{\mathrm{reg}}},\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}})^{\vee\vee}$ is a $(\omega_{0},\cdots,\omega_{n-2})$-stable reflexive Higgs sheaf with (1.3); Building on the stable case that has been clarified, $\mathcal{Q}_{X}^{\vee\vee}$ is locally free and its all Chern classes vanish.
  • (iii)
    $\mathcal{Q}_{X}^{\vee\vee}=\mathcal{Q}_{X}$ in codimension 22 (see Proposition 2.15); Buiding on Lemma 5.1, $\mathcal{Q}_{X}$ and $\mathcal{E}_{X}$ is locally free. In particular, the Higgs fields $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$ and $\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}}$ extend to the Higgs fields on $\mathcal{E}_{X}$ and $\mathcal{Q}_{X}$, respectively.
Summarizing the above statements and the induction hypothesis, the proof is complete. ∎
5.1.2. Independent of the choice of the polarization
Let us complete the remaining of Theorem 1.12 by proving the following.
Theorem 5.5.
Let $X$ be a compact Kähler klt space of dimension $n$, $\pi:\widehat{X}\rightarrow X$ be a resolution of singularities of $X$ and $(\mathcal{E}_{X},\theta_{X})$ be a locally free Higgs sheaf over $X$. Denote $(\mathcal{E}_{\widehat{X}},\theta_{\widehat{X}})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$. Then the following are equivalent:
  1. (i)
    There exists some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ such that $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with (1.3).
  2. (ii)
[…]
    ${\mathrm{ch}}_{1}(\mathcal{E}_{\widehat{X}})\cdot[\alpha_{0}]\cdots[\alpha_{n-2}]=0\ \text{and}\ {\mathrm{ch}}_{2}(\mathcal{E}_{\widehat{X}})=0.$
  3. (iii)
[…]
  4. (iv)
    All Chern classes of $\mathcal{E}_{X}$ vanish and $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$.
Ditto for polystability.
On the other hand, using Theorem 5.5 together with Theorem 5.4, Lemmas 2.21 and 2.24, one can easily verify that the definitions of the categories $\mathrm{Higgs}_{X},\mathrm{pHiggs}_{X},\mathrm{Higgs}_{X_{\mathrm{reg}}}$ and $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ are independent of the choice of the polarization $(\omega_{0},\cdots,\omega_{n-2})$.
Proof.
The implication $(iv)\Rightarrow(i)$ is obvious and $(iii)\Rightarrow(iv)$ follows from the item (4.1.1). Let us first verify $(i)\Rightarrow(ii)$. Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$, $\widehat{Y}$ be a desingularisation of the (unique) irreducible component of the fibre product $\widehat{X}\times_{X}Y$ and then we have the following commuting diagram
[…]
(5.4) $c_{1}(\mathcal{Q}_{k})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m}]\cdots[\alpha_{n-2}]=0.$
[…]
(5.5) $\begin{split}&\left(\mathrm{ch}_{2}(\mathcal{Q}_{k}^{\vee\vee})-\frac{1}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}}\cdot c_{1}(\mathcal{Q}_{k}^{\vee\vee})^{2}\right)\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\end{split}$
by letting $\delta\rightarrow 0$. Combining Proposition 2.15, the induction hypothesis (5.5) and the Hodge index Theorem (Lemma 5.3), it holds that
(5.6) $\begin{split}0&=\mathrm{ch}_{2}(\mathcal{E}_{X})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}2\cdot\mathrm{ch}_{2}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}\frac{1}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}}\cdot c_{1}(\mathcal{Q}_{k}^{\vee\vee})^{2}\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}\frac{\big(c_{1}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m}]\cdots[\alpha_{n-2}]\big)^{2}}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}\cdot[\pi^{*}\alpha_{m}]^{2}\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]}=0.\end{split}$
[…]
The proof of the polystable case is similar. ∎
Combining Theorem 5.5 with the classical non-abelian Hodge correspondence for a single Kähler class [17, 97, 86, 23] (see also [93, Theorem 1.9] for projective flat bundles), we immediately obtain its extension to multiple Kähler polarization on compact Kähler manifolds, which has already been established in the polystable case by [11].
Theorem 5.6 (Non-abelian Hodge theory for compact Kähler manifolds).
Let $X$ be a compact Kähler manifold. There exists an equivalence of categories $\mu_{X}:\mathrm{Higgs}_{X}\rightarrow\mathrm{LSys}_{X}.$ This equivalence restricts to an equivalence between $\mathrm{pHiggs}_{X}$ and $\mathrm{sLSys}_{X}$.
Remark 5.7.
This equivalence of categories in this theorem is indeed functorial with respect to the morphisms between compact Kähler manifolds, which can be easily seen from the construction of $\mu_{X}$ (see e.g. [93, Section 9]).
5.2. Proof of Theorem 1.15
Now let us show Theorem 1.15. Suppose that $f:Z\rightarrow X$ is a bimeromorphic, holomorphic map between compact Kähler klt spaces and $(\mathcal{E}_{Z},\theta_{Z})\in\mathrm{Higgs}_{Z}$, we need to construct a $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ such that $\pi^{*}(\mathcal{E}_{X},\theta_{X})=(\mathcal{E}_{Z},\theta_{Z})$.
Let $g:W\rightarrow X$ be an orbifold modification of $X$ and $Y$ be a desingularisation of the (unique) irreducible component of the fibre product $W\times_{X}Z$. Then we have the following commuting diagram
[…]
In particular, $\pi_{1}:Y\rightarrow W$, $\pi_{2}:Y\rightarrow Z$ and $\pi:Y\rightarrow X$ are resolution of singularities of $Z$, $W$ and $X$, respectively. Recall $(\mathcal{E}_{Y},\theta_{Y}):=\pi_{2}^{*}(\mathcal{E}_{Z},\theta_{Z})\in\mathrm{Higgs}_{Y}$ by Theorem 5.5. It reduces to show the following statement by Lemma (4.1.1).
Proposition 5.8.
There exists some locally free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ on $X$ such that $(\mathcal{E}_{Y},\theta_{Y})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$
Let us set the stage with some preparations. Suppose that $E_{Y}=\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{LSys}_{Y}$ and define $E_{X}=(\pi^{*})^{-1}E_{Y}\in\mathrm{LSys}_{X}$. Denote by $E_{\rho_{X}}$ and $E_{\rho_{Y}}$ the corresponding local systems.
Suppose that that $E_{X}$ is induced by a linear representation $\rho_{X}:\pi_{1}(X)\rightarrow\mathrm{GL}(r,\mathbb{C})$ in block upper-triangular form
$\begin{pmatrix}\rho_{1}&*&\cdots&*\\ 0&\rho_{2}&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\rho_{m}\end{pmatrix}$,
where each diagonal block $\rho_{i}:\pi_{1}(X)\to\mathrm{GL}(r_{i},\mathbb{C})$ is irreducible. This induces an increasing filtration of local systems
$0=E_{0}\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{m}=E_{X},$
where each $E_{i}$ corresponds to the truncated representation
$\widehat{\rho}_{i}=\begin{pmatrix}\rho_{1}&*&\cdots&*\\ 0&\rho_{2}&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\rho_{i}\end{pmatrix}$.
In particular, the quotient $Q_{X,i}:=E_{X,i}/E_{X,i-1}$ is a simple local system induced by $\rho_{i}$. Then $E_{W}=g^{*}E_{X}$ admits a filtration of local systems
(5.8) $0\subsetneq E_{W,1}\subsetneq\cdots\subsetneq E_{W,m}=E_{W}$
such that each $E_{W,l}=g^{*}E_{X,l}$ is induced by $\widehat{\rho_{i}}\circ g_{*}$. Because $g_{*}:\pi_{1}(W)\rightarrow\pi_{1}(X)$ is an isomorphism by Takayama’s result [101], the quotient
$Q_{W,i}=E_{W,i}/E_{W,i-1}=g^{*}Q_{X,i}$
is a simple local system induced by $\rho_{i}\circ g_{*}$. Analogously, the flat bundle $(V_{Y},D_{Y})$ induced by $E_{Y}$ admits a filtration of flat bundles
(5.9) $0\subsetneq V_{Y,1}\subsetneq\cdots\subsetneq V_{Y,m}=V_{Y}$
such that each $V_{Y,l}$ is $D$-flat and induced by $\widehat{\rho_{i}}\circ\pi_{*}$, and the quotient $Q_{Y,i}=V_{Y,i}/V_{Y,i-1}$ with the induced flat connection $D_{Y,i}$ is induced by $\pi^{*}\rho_{i}\in\mathrm{sRep}_{Y}$.
By applying non-abelian Hodge correspondence for compact Kähler manifolds in [23], $(Q_{Y,i},\theta_{Y,i})$ admits a harmonic metric $h_{i}$. Recall there is a unique harmonic decomposition
$D_{Y,i}=\partial_{Y,i}+\bar{\partial}_{Y,i}+\theta_{Y,i}+\theta_{Y,i}^{*}$
and $Q_{Y_{i}}$ can be associate with a Higgs bundle structure $(\mathcal{Q}_{Y,i},\theta_{Y,i})=(Q_{Y,i},\bar{\partial}_{Y,i},\theta_{Y,i}).$ Building on Proposition 4.11 and Theorem 1.12, there exists a locally Higgs sheaf $(\mathcal{Q}_{X,i},\theta_{X,i})\in\mathrm{pHiggs}_{X}$ on $X$ such that $(\mathcal{Q}_{Y,i},\theta_{Y,i})=\pi^{*}(\mathcal{Q}_{X,i},\theta_{X,i}).$ According to the above construction, the underlying $C^{\infty}$-bundle of $(\mathcal{Q}_{X,i},\theta_{X,i})$ is isomorphic to the underlying $C^{\infty}$-bundle of $Q_{X,i}$. In particular,
$(\mathcal{Q}_{W,i},\theta_{W,i}):=g^{*}(\mathcal{Q}_{X,i},\theta_{X,i})\in\mathrm{pHiggs}_{W}$
and its underlying $C^{\infty}$-bundle is isomorphic to the underlying $C^{\infty}$-bundle of $Q_{W,i}$ ($C^{\infty}$ makes sense via local embeddings of analytic spaces).
Now let us complete the proof of Proposition 5.8. The basic idea is constructing a locally free Higgs sheaf $(\mathcal{E}_{W},\theta_{W})$ of the underlying $C^{\infty}$-bundle of $E_{\rho_{W}}$ via the filtration $\{E_{W,k}\}_{k=1}^{m}$ over a standard orbifold structure $W_{\mathrm{orb}}=\{U_{\alpha},G_{\alpha},\mu_{\alpha}\}$ of $W$ and showing $(g_{*}\mathcal{E}_{W})^{\vee\vee}$ is locally free.
Proof of Proposition 5.8.
Let $(V_{\mathrm{orb}},D_{\mathrm{orb}})$ be the flat orbi-bundle induced by $\{\mu_{\alpha}^{*}E_{\rho_{W}}\}$. It admits a filtration of flat orbi-bundles
(5.10) $0=V_{\mathrm{orb},0}\subsetneq V_{\mathrm{orb},1}\subsetneq\cdots\subsetneq V_{\mathrm{orb},m}=V_{\mathrm{orb}}$
such that each $V_{\mathrm{orb},i}$ in invariant under $D_{\mathrm{orb}}$ and the quotient $Q_{\mathrm{orb},i}=V_{\mathrm{orb},i}/V_{\mathrm{orb},i}$ with the induced flat connection $D_{\mathrm{orb},i}$ is induced by $\{\mu_{\alpha}^{*}Q_{W_{i}}\}$. Note that $Q_{\mathrm{orb},i}$ has a Higgs orbi-bundle structure $\{Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i}\}$ induced by $\{\mu_{\alpha}^{*}(\mathcal{Q}_{W,i},\theta_{W,i})\}$. Then $\{Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i}\}$ is polystable with respect to some orbifold Kähler class $\omega_{\mathrm{orb}}$ by Theorem 5.5 and [110, Proposition 4.8], and all orbifold Chern classes vanish because $Q_{\mathrm{orb},i}$ is flat. Building on the existence of the Hermitian-Einstein metric of stable Higgs orbi-bundles due to [110], $(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})$ admits an orbifold pluri-harmonic metric $H_{\mathrm{orb},i}$. Since $\mu_{\alpha}$ is étale over $W_{\mathrm{reg}}$, then $H_{\mathrm{orb},i}$ descend to a pluri-harmonic metric $H_{i}$ of $(\mathcal{Q}_{W,i},\theta_{W,i})|_{W_{\mathrm{reg}}}$. By virtue of the uniqueness by Theorem 1.13 (1), $D_{\theta_{W,i},H_{i}}$ identifies with the flat connection structure of $Q_{W,i}|_{W_{\mathrm{reg}}}$ and so $D_{\theta_{\mathrm{orb},i},H_{\mathrm{orb},i}}=D_{\mathrm{orb},i}.$
Note that the Hodge isomorphism Theorem holds in the context of compact complex orbifolds (cf. [77, Proposition 2.1]), so we could completely follow the argument of [23, Proof of Theorem A] to construct a Higgs orbi-bundle structure $(V_{\mathrm{orb}},\bar{\partial}_{\mathrm{orb}},\theta_{\mathrm{orb}})$ of $V_{\mathrm{orb}}$ such that each $V_{\mathrm{orb},i}$ is $\bar{\partial}_{\mathrm{orb}}$-invariant and $\theta_{\mathrm{orb}}$-invariant, and the induced Higgs bundle structures on $Q_{\mathrm{orb},i}$ coincides with $(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})$. Consequently,
$(\mathcal{E}_{\mathrm{orb}},\theta_{\mathrm{orb}}):=(V_{\mathrm{orb}},\bar{\partial}_{\mathrm{orb}},\theta_{\mathrm{orb}})$
is the extension of locally Higgs sheaf $\{\mu_{\alpha}^{*}(\mathcal{Q}_{W,i},\theta_{W,i})\}_{i=1}^{m}$. By applying [33, Lemma A.4], $\left((\mu_{\alpha})_{*}(\mu_{\alpha}^{*}\mathcal{Q}_{W,i})\right)^{G_{\alpha}}$ is reflexive and coincides with the locally free sheaf $\mathcal{Q}_{W,i}$ since $\mu_{\alpha}$ is étale over $W_{\mathrm{reg}}$; $\mathcal{E}_{W}:=\left((\mu_{\alpha})_{*}\mathcal{E}_{\alpha}\right)^{G_{\alpha}}$ is reflexive and admits a filtration of saturated subsheaves
$0=\mathcal{E}_{W,0}\subsetneq\mathcal{E}_{W,1}\subsetneq\cdots\subsetneq\mathcal{E}_{W,m}=\mathcal{E}_{W}$
such that $\mathcal{E}_{W,i}/\mathcal{E}_{W,i-1}=\mathcal{Q}_{W,i}$, where $\mathcal{E}_{W,l}=\left((\mu_{\alpha})_{*}(\mathcal{E}_{\mathrm{orb},i})_{\alpha})\right)^{G_{\alpha}}$ and $\mathcal{E}_{\mathrm{orb},i}=(V_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb}}|_{V_{\mathrm{orb}},i})$. Lemma 5.1 implies that each $\mathcal{E}_{W,i}$ is locally free and so is $\mathcal{E}_{W}$. Moreover, the orbifold Higgs field $\theta_{\mathrm{orb}}$ descends first to a Higgs field on $\mathcal{E}_{W}|_{W_{\mathrm{reg}}}$, and then extends to a Higgs field $\theta_{W}$ on $\mathcal{E}_{W}$ by Lemma 2.4, which gives a locally free Higgs sheaf structure of the underlying $C^{\infty}$-bundle of $E_{\rho_{W}}$. Then $(V_{Y},\bar{\partial}_{Y},\theta_{Y})=\pi_{1}^{*}(\mathcal{E}_{W},\theta_{W})$ is a locally free Higgs sheaf structure of $(V_{Y},D_{Y})$.
From the above construction, each $V_{Y,i}$ is $\theta_{Y}$-invariant and $\bar{\partial}_{Y}$-invariant, and the induced Higgs structure of $(V_{Y},\bar{\partial}_{Y},\theta_{Y})$ on the quotient $Q_{Y,i}$ coincides with $(Q_{Y,i},\bar{\partial}_{Y,i},\theta_{Y,i})$ (because these hold on a Zariski open subset of $Y$). According to the explicit construction of the non-abelian Hodge correspondence for compact Kähler manifolds (cf. [93, Section 9] and [23]), we have $\mu_{Y}(V_{Y},\bar{\partial}_{Y},\theta_{Y})={E_{Y}}=\mu_{Y}(\mathcal{E}_{Y},\theta_{Y}),$ which implies $(\mathcal{E}_{X},\theta_{X})=\pi_{1}^{*}(\mathcal{E}_{W},\theta_{W})$.
Finally, let us verify that the reflexive sheaf $\mathcal{E}_{X}:=(g_{*}\mathcal{E}_{W})^{\vee\vee}$ is locally free, which completes the proof. $\theta_{W}$ induces a Higgs field of $\mathcal{E}_{X}$ in codimension $2$ and so induces a Higgs field $\theta_{X_{\mathrm{reg}}}$ of $\mathcal{E}_{X_{\mathrm{reg}}}$. By the above construction, $\mathcal{E}_{X}$ admits a filtration of reflexive sheaves
$0=\mathcal{E}_{X,0}\subsetneq\mathcal{E}_{X,1}\subsetneq\cdots\subsetneq\mathcal{E}_{X,m}=\mathcal{E}_{X},$
where $\mathcal{E}_{X,i}:=(g_{*}\mathcal{E}_{W,i})^{\vee\vee}$. It is straightforward to see that each $\mathcal{E}_{X,i}$ is saturated in $\mathcal{E}_{X}$. This follows from the reflexivity of these sheaves and from the fact that $\mathcal{E}_{X,i}/\mathcal{E}_{X,i-1}=\mathcal{Q}_{X,i}$ outside the indeterminacy locus $S$ of $g^{-1}$. Again by reflexivity, we have
$(\mathcal{E}_{X,i}/\mathcal{E}_{X,i-1})^{\vee\vee}=\mathcal{Q}_{X,i},$
and its orbifold Chern classes vanish. Recall that $S$ has codimension at least $3$; consequently, the orbifold Chern classes of each $\mathcal{E}_{X,i}$ vanish because $g_{\mathrm{orb}}^{[*]}\mathcal{E}_{X,i}$ coincides with $E_{\mathrm{orb},i}$ outside $g^{-1}(S)$ and all orbifold Chern classes of $E_{\mathrm{orb},i}$ vanish. Combining these statements with Proposition 2.15, we deduce that $\mathcal{E}_{X,i}/\mathcal{E}_{X,i-1}$ is locally free in codimension $3$. Applying Lemma 5.1 then shows that each $\mathcal{E}_{X,i}$ is locally free by induction on $i=1,\cdots,m$, and hence so is $\mathcal{E}_{X}$. ∎
5.3. Construction of the correspondence
Now let us complete the proof of Theorem 1.3 and Theorem 1.7. We will always assume that $X$ is a compact Kähler klt space.
5.3.1. Proof of Theorem 1.3
We begin with constructing the functor $\mu_{X}$. For any $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$, Theorem 5.5 implies $\pi^{*}(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{\widehat{X}}$ for a resolution of singularities $\pi:\widehat{X}\rightarrow X$. Combining the item (4.1.2), we could define $\mu_{X}$ as below.
Definition 5.9.
$\mu_{X}:\mathrm{Higgs}_{X}\rightarrow\mathrm{LSys}_{X}$ is defined by
$(\mathcal{E}_{X},\theta_{X})\mapsto(\pi^{*})^{-1}\left(\mu_{\widehat{X}}(\pi^{*}(\mathcal{E}_{X},\theta_{X}))\right),$
where $\pi$ is a resolution of singularities $\pi:\widehat{X}\rightarrow X$.
It can be easily seen that the diagram (1.2) commutes. It’s necessary to check that $\mu_{X}$ is independent of the choice of $\pi$. Let $\pi_{1}:X_{1}\rightarrow X$ and $\pi_{2}:X_{2}\rightarrow X$ be two resolution of singularities, we need to verify the following statement.
Lemma 5.10 (The well-definedness of $\mu_{X}$).
$(\pi_{1}^{*})^{-1}\left(\mu_{X_{1}}(\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X}))\right)=(\pi_{2}^{*})^{-1}\left(\mu_{X_{2}}(\pi_{2}^{*}(\mathcal{E}_{X},\theta_{X}))\right)$
Proof.
Let $\widehat{X}$ be a desingularisation of the (unique) irreducible component of the fibre product $X_{1}\times_{X}X_{2}$ and so we have the following natural commuting diagram
[…]
Note that
$g^{*}(\pi_{1}^{*})^{-1}\left(\mu_{X_{1}}(\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X}))\right)$ $=f_{1}^{*}\left(\mu_{X_{1}}(\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X}))\right)=\mu_{\widehat{X}}\left(f_{1}^{*}\pi_{1}^{*}(\mathcal{E}_{X},\theta_{X})\right)$
$=\mu_{\widehat{X}}\left(f_{2}^{*}\pi_{2}^{*}(\mathcal{E}_{X},\theta_{X})\right)=f_{2}^{*}\left(\mu_{X_{2}}(\pi_{2}^{*}(\mathcal{E}_{X},\theta_{X}))\right)$
$=g^{*}(\pi_{2}^{*})^{-1}\left(\mu_{X_{2}}(\pi_{2}^{*}(\mathcal{E}_{X},\theta_{X}))\right)$
where the second equality and the fourth equality follow from the functoriality of the non-abelian Hodge correspondence for compact Kähler manifolds. Because $g$ is also a resolution of singularities and so $g^{*}$ is an isomorphism by the item (4.1.2), which completes the proof. ∎
According to the above construction, it’s obvious that if $(\mathcal{E}_{X},\theta_{X})\in\mathrm{pHiggs}_{X}$ then $\mu_{X}(\mathcal{E}_{X},\theta_{X})\in\mathrm{sLSys}_{X}$. Also, we have the following functoriality building on the existence of functorial resolution of singularities (see e.g. [22, Theorem 3.10]) and the functoriality of $\mu_{X}$ for compact Kähler manifolds.
Corollary 5.11.
$\mu_{X}$ is functorial with respect to local analytic isomorphisms.
To prove that $\mu_{X}$ is one-to-one, it suffices to construct its inverse map. Let $E_{X}\in\mathrm{LSys}_{X}$. Suppose that $\pi^{*}E_{X}\in\mathrm{LSys}_{\widehat{X}}$. It follows from Theorem 1.15 that there exists $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ such that $\mu_{\widehat{X}}^{-1}(\pi^{*}E_{X})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$.
Definition 5.12.
$\eta_{X}:\mathrm{LSys}_{X}\rightarrow\mathrm{Higgs}_{X}$ is defined by sending $E_{X}$ to $(\mathcal{E}_{X},\theta_{X})$ which is constructed as above.
It follows immediately from the above construction that $\eta_{X}\circ\mu_{X}=\operatorname{Id}$, $\mu_{X}\circ\eta_{X}=\operatorname{Id}$ and $\mu_{X}$ satisfies the properties list in Theorem 1.3.
Remark 5.13.
One could verify that $\mu_{X}$ is in fact functorial with respect to any morphism between compact Kähler klt spaces as in [38, page 318].
5.3.2. Proof of Theorem 1.7
Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$ induced by some $G\subset\mathrm{Aut}(Y)$. For any $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$, $(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})=\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is locally free and uniquely extends to a locally Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{Higgs}_{Y}$ by Theorem 1.12. This also implies $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free because $\gamma$ is étale over $X_{\mathrm{reg}}$. Then $\gamma^{*}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\rightarrow\mathrm{Higgs}_{Y}$ is defined by sending $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ to $(\mathcal{E}_{Y},\theta_{Y})$.
By the reflexivity, $(\mathcal{E}_{Y},\theta_{Y})$ is $G$-equivariant and thus $\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{LSys}_{X}$ is $G$-equivariant by Corollary 5.11. Recall $\gamma$ is finite and étale over $X_{\mathrm{reg}}$, $\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})|_{\gamma^{-1}(X_{\mathrm{reg}})}$ descents to a local system $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$.
Definition 5.14.
$\mu_{X_{\mathrm{reg}}}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\rightarrow\mathrm{LSys}_{X_{\mathrm{reg}}}$ is defined by
$(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\mapsto\left(\gamma^{\circ}_{*}(\mu_{Y}(\gamma^{*}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})))\right)^{G},$
where $\gamma:Y\rightarrow X$ is a maximally quasi-étale cover of $X$ and $\gamma^{\circ}=\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})}$.
To illustrate that $\mu_{X_{\mathrm{reg}}}$ is well-defined, it’s necessary to check that
Lemma 5.15.
$\mu_{X_{\mathrm{reg}}}$ is independent of the choice of maximally quasi-étale covers.
Proof of Lemma 5.15.
Let $\gamma_{1}:Y_{1}\rightarrow X$ and $\gamma_{2}:Y_{2}\rightarrow X$ be two maximally quasi-étale covers of $X$. Then the fiber product $Y:=Y_{1}\times_{X}Y_{2}$ is a quasi-étale cover of $Y_{1}$. By Definition 4.2 (1), $Y$ is a étale cover of $Y_{1}$ and thus is a maximally quasi-étale cover of $X$. Then, to prove the statement, we may assume that $\gamma_{12}:Y_{1}\rightarrow Y_{2}$ is an étale cover. We adopt the associated symbols as in the aforementioned construction. By the construction. it suffices to check that $E_{Y_{1,\mathrm{reg}}}=\gamma_{12}^{*}E_{Y_{2,\mathrm{reg}}}$. Due to the reflexivity, we in fact have $(\mathcal{E}_{Y_{1}},\theta_{Y_{1}})=\gamma_{12}^{*}(\mathcal{E}_{Y_{2}},\theta_{Y_{2}})$. Applying Corollary 5.11, the proof is complete. ∎
To verify that $\mu_{X_{\mathrm{reg}}}$ is one-to-one, it suffices to construct its inverse map. Let $E_{X_{\mathrm{reg}}}\in\mathrm{LSys}_{X_{\mathrm{reg}}}$. The item (4.6.4) implies that $(\gamma|_{\gamma^{-1}(X_{\mathrm{reg}})})^{*}E_{X_{\mathrm{reg}}}$ uniquely extends to some $G$-equivariant local system $E_{Y}\in\mathrm{LSys}_{Y}$. Then $\gamma^{*}:\mathrm{LSys}_{X_{\mathrm{reg}}}\rightarrow\mathrm{LSys}_{Y}$ is defined by sending $E_{X_{\mathrm{reg}}}$ to $E_{Y}$. Building on Corollary 5.11 and the fact that $\mu_{Y}$ is one-to-one, $\mu_{Y}^{-1}(E_{Y})\in\mathrm{Higgs}_{Y}$ is $G$-equivariant and thus $\mu_{Y}^{-1}(E_{Y})|_{\gamma^{-1}(X_{\mathrm{reg}})}$ descend to some $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$ (again by the item ((4.6.3))).
Definition 5.16.
$\eta_{X_{\mathrm{reg}}}:\mathrm{LSys}_{X_{\mathrm{reg}}}\rightarrow\mathrm{Higgs}_{X_{\mathrm{reg}}}$ is defined by
$E_{X_{\mathrm{reg}}}\mapsto\left(\gamma^{\circ}_{*}(\mu_{Y}^{-1}(\gamma^{*}E_{X_{\mathrm{reg}}}))\right)^{G}$
It follows immediately from the construction that $\mu_{X_{\mathrm{reg}}}\circ\eta_{X_{\mathrm{reg}}}=\operatorname{Id}$, $\eta_{X_{\mathrm{reg}}}\circ\mu_{X_{\mathrm{reg}}}=\operatorname{Id}$ and $\mu_{X_{\mathrm{reg}}}$ satisfies the remaining part of Theorem 1.7. The proof is complete.
5.4. Quasi-uniformization results
In this subsection, we prove Corollary 1.9 and Corollary 1.10 and Theorem 1.11.
Proof of Corollary 1.9.
Let $X$ be a compact Kähler klt space of dimension $n$ with $K_{X}$ numerically trivial. Since $\mathcal{T}_{X}$ is generically nef with respect to Kähler form $\omega_{0},\cdots,\omega_{n-2}$ by [53, Proposition 6.5], $\mathcal{T}_{X}$ is semi-stable with respect to any Kähler classes $\omega_{0},\cdots,\omega_{n-2}$ since $c_{1}(K_{X})=0$. Then (1.5) follows from the Bogomolov inequality (3.1).
Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover. Note that $\mathcal{T}_{Y}=\gamma^{[*]}\mathcal{T}_{X}$ by the reflexivity. Since $\widehat{c}_{1}(\mathcal{T}_{X})=0$, it follows from Theorem 1.12 that the equality in (1.5) holds if and only if $\mathcal{T}_{Y}$ is a locally free and flat sheaf, which is equivalent to that $Y$ is smooth by the Lipman-Zariski conjecture for klt spaces (cf. [63, Theorem 1.13]) and so is a torus thanks to Yau’s theorem [109]. The proof is complete. ∎
Proof of Corollary 1.10.
The proof is similar to [55, Proposition 2.20]. Let $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{{X_{\mathrm{reg}}}})$ be a semi-stable reflexive Higgs sheaves. Then [110, Corollary 1.3] implies that $\mathcal{G}_{X_{\mathrm{reg}}}:=\operatorname{End}(\mathcal{E}_{X_{\mathrm{reg}}})$ with the induced Higgs field $\theta_{\mathcal{G}_{X_{\mathrm{reg}}}}$ is semi-stable. The inequality (1.6) has shown in Theorem 3.2. Recall (3.13) that $\mathcal{E}_{X}$ satisfies the equality in (1.3) if and only if $\mathcal{G}_{X}$ satisfies $\widehat{c}_{1}(\mathcal{G}_{X})=0$ and (1.3). Building on Theorem 1.12, we complete the proof. ∎
Let us outline the proof of Theorem 1.11 by building on the results in Sections 3 and 4, following the spirits of [53] and [58, Section 4]. We refer to [53, Section 3.2 and 4.2] for the precise definitions of stability and of the intersection number taken with respect to the non‑pluripolar product.
Proof of Theorem 1.11.
Let $X$ be a projective klt variety of dimension $n$ with $K_{X}$ big. There is a natural Higgs sheaf on $X$ consists of the reflexive sheaf $\mathcal{E}_{X}:=\Omega_{X}^{[1]}\oplus\mathcal{O}_{X}$ and the Higgs field $\theta_{X}$ defined by
(5.12) $\theta_{X}:\Omega_{X}^{[1]}\oplus\mathcal{O}_{X}\longrightarrow(\Omega_{X}^{[1]}\oplus\mathcal{O}_{X})\otimes\Omega_{X}^{[1]},\quad a\oplus b\longmapsto(0\oplus 1)\otimes a.$
Since $K_{X}$ is big, $X$ admits a canonical model $X_{\text{can}}$ by [4, Theorem 1.2] and denote $h:X\dashrightarrow X_{\text{can}}$ be the birational map induced by the $K_{X}$-MMP. Let $p_{1}:X_{1}\rightarrow X$ and $q_{1}:X_{1}\rightarrow X_{\text{can}}$ be resolutions of $f$. Then there exists an $q_{1}$-effective divisor $E$ such that
$p_{1}^{*}K_{X}=q_{1}^{*}K_{\text{can}}+E$
and $p_{1}(\mathrm{supp}(E))$ coincides with the $h$-exceptional divisor. By [21, Lemma A.5] (see also [53, Proposition 3.15]), we have
(5.13) $\langle p_{1}^{*}K_{X}^{k}\rangle=q_{1}^{*}K_{X_{\text{can}}}^{k},\ \forall k\in\mathbb{N}.$
Let us recall the definition of $\widehat{\Delta}(\mathcal{E}_{X})\cdot\langle K_{X}^{n-2}\rangle$ introduced in [53, Section 4.2.1]. We take an orbifold modification $f^{\prime}:Y^{\prime}\rightarrow X$ such that $Y^{\prime}$ admits a standard orbifold structure $Y_{\mathrm{orb}}^{\prime}:=\{U^{\prime}_{\mu},G_{\mu}^{\prime},\pi_{\mu}^{\prime}\}$. $\mathcal{E}_{Y^{\prime}}:=f^{\prime*}\mathcal{E}_{X}/\mathrm{tor}$ induces a torsion-free orbi-sheaf $\mathcal{E}_{\mathrm{orb}}^{\prime}:=\{\widehat{\pi}_{\mu}^{*}\mathcal{F}_{Y^{\prime}}/\mathrm{tor}\}$. According to Lemma 2.2, there exists a projective bimeromorphic morphism $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow Y_{\mathrm{orb}}^{\prime}$ from an effective complex orbifold $Y_{\mathrm{orb}}=\{U_{\mu},G_{\mu},\pi_{\mu}\}$ to $Y_{\mathrm{orb}}^{\prime}$ such that the torsion-free pullback of $\mathcal{E}_{\mathrm{orb}}^{\prime}$ is an orbi-bundle $E_{\mathrm{orb}}^{\prime}$. Set $f:=f^{\prime}\circ g$ and it follows from [53, Proposition 4.18]
(5.14) $\left(2\widehat{c}_{2}(X)-\frac{n}{n+1}\widehat{c}_{1}^{2}(X)\right)\cdot\langle K_{X}^{n-2}\rangle=\left(2c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}}^{\prime})-\frac{n}{n+1}({c}_{1}^{\mathrm{orb}})^{2}(E_{\mathrm{orb}}^{\prime})\right)\cdot\langle f^{*}K_{X}^{n-2}\rangle.$
Note that on every orbifold chart $U_{\mu}$, $(f\circ\pi_{\mu})^{*}\mathcal{E}_{X}$ admits a functorial Higgs field $(f\circ\pi_{\mu})^{*}\theta_{X}$ by [63, Theorem 1.10] (cf. [37, Construction 5.5]), which induces a Higgs field $\theta_{\mathrm{orb}}^{\prime}$ of $E_{\mathrm{orb}}^{\prime}$ because $E_{\mathrm{orb}}^{\prime}=\{(f\circ\pi_{\mu})^{*}\mathcal{E}_{X}/\mathrm{tor}\}$.
Choose resolutions $p_{2}:X_{2}\to Y$ and $q_{2}:X_{2}\to X_{1}$ of the composition $p^{-1}\circ f:Y\dashrightarrow X_{1}$. Let $V_{\mu}$ be a functorial resolution of singularities of the normalization of $U_{\mu}\times_{Y}X_{2}$; then the group $G_{\mu}$ lifts to a subgroup $\widehat{G}_{\mu}\subset\operatorname{Aut}(V_{\mu})$. Denote by $\widehat{\pi}_{\mu}:V_{\mu}\to V_{\mu}/\widehat{G}_{\mu}$ the quotient map. The triple $\{V_{\mu},\widehat{G}_{\mu},\widehat{\pi}_{\mu}\}$ forms a complex effective orbifold $Z_{\mathrm{orb}}$; we write $Z$ for its underlying complex space, $g_{1}:Z\rightarrow Y$ and $g_{2}:Z\rightarrow X$ for the induced bimeromorphisms. All these morphisms are summarised in the following diagram:
[…]
Denote $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ be the pull-back of the Higgs orbi-bundle $(E_{\mathrm{orb}}^{\prime},\theta_{\mathrm{orb}}^{\prime})$ by $g_{1}$. Note that (5.13) implies
(5.15) $\langle g_{1}^{*}f^{*}K_{X}^{k}\rangle=\langle g_{2}^{*}q_{2}^{*}p_{1}^{*}K_{X}^{k}\rangle=g_{2}^{*}q_{2}^{*}q_{1}^{*}K_{X}^{k}.$
Set $p:=f\circ g_{1}:Z\rightarrow X$ and $q:=q_{1}\circ q_{2}\circ g_{2}:Z\rightarrow X_{\text{can}}$. Using the same argument of the proof of [53, Lemma 4.14 (1)], we have that $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ is $\langle g_{1}^{*}f^{*}K_{X}^{n-1}\rangle$-stable and so is $q^{*}K_{\text{can}}^{n-1}$-stable. Furthermore, the Poincare duality holds for orbifolds, we have that $\langle g_{1}^{*}f^{*}K_{X}^{k}\rangle=\pi_{*}\langle\pi^{*}g_{1}^{*}f^{*}K_{X}^{k}\rangle=g_{1}^{*}\langle f^{*}K_{X}^{k}\rangle$ in $H^{2n-4}(Z,\mathbb{R})$ according to the definition of the non-pluripolar product [53, Section 3.1.3], then (5.15) and (5.14) yield
(5.16) $\left(2\widehat{c}_{2}(X)-\frac{n}{n+1}\widehat{c}_{1}^{2}(X)\right)\cdot\langle K_{X}^{n-2}\rangle=\left(2c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})-\frac{n}{n+1}({c}_{1}^{\mathrm{orb}})^{2}(E_{\mathrm{orb}})\right)\cdot q^{*}K_{X}^{n-2}.$
Recall [40, Theorem A] that the reflexive cotangent sheaf $\Omega_{X_{\text{can}}}^{[1]}$ is $K_{X_{\text{can}}}^{n-1}$-semistable, which implies that the natural Higgs sheaf $(\mathcal{E}_{X_{\text{can}}},\theta_{X_{\text{can}}})$ on $X_{\text{can}}$ is $K_{X_{\text{can}}}^{n-1}$-stable (cf. [55, Proposition 2.8]). According to the above construction, $q:Z\rightarrow X_{\text{can}}$ satisfies all properties listed in Proposition 3.4 by substituting $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$= $(\mathcal{E}_{X_{\text{can}}},\theta_{X_{\text{can}}})|_{(X_{\text{can}})_{\mathrm{reg}}}$ and $\omega_{0}=\cdots=\omega_{n-2}=\omega$, where $\omega$ is a Kähler form in the class $K_{X}$. Under the assumption that the equality in (1.8) holds, applying Theorem 3.5 and a similar computation of Section 3.1.3 give
(5.17) $0=4\pi^{2}\left(2c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})-\frac{n}{n+1}({c}_{1}^{\mathrm{orb}})^{2}(E_{\mathrm{orb}})\right)\cdot q^{*}K_{X}^{n-2}\geq\int_{X_{\text{can}}\setminus\Sigma}\|F_{H_{\infty},\theta_{\text{can}}}^{\perp}\|^{2}_{H_{\infty},\omega}\cdot\frac{\omega^{n}}{n!}$
and so $F_{H_{\infty},\theta_{\text{can}}}^{\perp}\equiv 0$ on $X\setminus Z$. In particular, $\mathcal{G}_{X_{\text{can}}}:=\operatorname{End}(\mathcal{E}_{X_{\text{can}}})$ with the induced Higgs field $\theta_{\mathcal{G}}$ admits a pluri-harmonic metric on $X_{\text{can}}\setminus\Sigma$ and so admits a pluri-harmonic metric on the regular locus of $X_{\text{can}}$ by Proposition 4.8. Then it follows from Proposition 4.14 that for a maximally quasi-étale cover $\gamma:Y\rightarrow X_{\text{can}}$, $\gamma^{[*]}\mathcal{G}_{X_{\text{can}}}$ is locally free and its all Chern classes vanish. By the reflexivity, $\gamma^{[*]}\mathcal{G}_{X_{\text{can}}}=\operatorname{End}(\Omega_{Y}^{[1]}\oplus\mathcal{O}_{Y})$. As $\mathcal{T}_{Y}$ is a direct summand of $\operatorname{End}(\Omega_{Y}^{[1]}\oplus\mathcal{O}_{Y})$, it follows that $\mathcal{T}_{Y}$ is locally free. Therefore $Y$ is smooth by the solution of Lipman-Zariski conjecture for klt spaces [33, 63] and $Y$ satisfies the Miyaoka-Yau equality with respect to $K_{Y}^{n-2}$. Since $K_{Y}=\gamma^{*}K_{X}$ is ample, the proof is complete by the classical result of Yau [108]. ∎
Remark 5.17.
Although [58, Theorem 1.2] assumes that $X$ is smooth in codimension $2$, this condition is in fact unnecessary. The general case follows by combining the argument above with the methods of Section 4 in the same reference.